### v1

# CLEME2.0: Towards More Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction

Abstract
The paper focuses on improving the interpretability of Grammatical Error Correction (GEC) metrics, which receives little attention in previous studies. To bridge the gap, we propose CLEME2.0, a reference-based evaluation strategy that can describe four elementary dimensions of GEC systems, namely hit-correction, error-correction, under-correction, and over-correction. They collectively contribute to revealing the critical characteristics and locating drawbacks of GEC systems. Evaluating systems by Combining these dimensions leads to high human consistency over other reference-based and reference-less metrics. Extensive experiments on 2 human judgement datasets and 6 reference datasets demonstrate the effectiveness and robustness of our method.111All the codes will be released after the peer review.
CLEME2.0: Towards More Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction
Jingheng Ye1††thanks: indicates equal contribution., Zishan Xu1∗, Yinghui Li1, Xuxin Cheng2, Linlin Song3, Qingyu Zhou, Hai-Tao Zheng1,4††thanks: Corresponding author: Hai-Tao Zheng. (E-mail: zheng.haitao@sz.tsinghua.edu.cn), Ying Shen5, Xin Su6 1Tsinghua University, 2Peking University 3Huazhong University of Science and Technology 4Peng Cheng Laboratory, 5Sun-Yat Sen University, 6Tencent {yejh22,xzs23}@mails.tsinghua.edu.cn
1 Introduction
Grammatical Error Correction (GEC) is the task of automatically detecting and correcting all grammatical errors in a given text Li et al. (2022a, b); Ma et al. (2022); Ye et al. (2022). A core component of any NLP tasks is the development of automatic metrics that can objectively measure model performance Bryant et al. (2023). However, proposing appropriate evaluation of GEC has long been a challenging task Madnani et al. (2011); Li et al. (2023a), due to the subjectivity Bryant and Ng (2015), complexity Mita et al. (2019) and subtlety Choshen and Abend (2018) of GEC Napoles et al. (2015).
Recent studies have been trying to develop GEC metrics that can achieve high correlations with human judgements Yoshimura et al. (2020a), with less attention paid to the interpretability of the automatic metrics. We define the interpretability of metrics as their ability to reveal the concerned characteristics of systems, which is vital in locating the drawbacks of a certain system. It is well-acknowledged that excellent GEC systems, which usually conform to the principle of minimal editing, should adhere to two gold principles, namely grammaticality and faithfulness. Grammaticality necessitates that all grammatical errors should be accurately corrected, while faithfulness requires that the corrections maintain the original textual meaning and syntactic structure. However, the widely-adopted mainstream GEC metrics Bryant et al. (2017); Ye et al. (2023c) indicate the GEC performance by precision, recall, and F scores, which can hardly characterize these critical dimensions of GEC systems, thus hindering the development.
Therefore, we propose CLEME2.0, a more interpretable reference-based evaluation strategy that can describe four fundamental aspects of GEC systems: hit-correction, error-correction, under-correction, and over-correction. The first three aspects are responsible for describing grammaticality, while the last one is for faithfulness since the over-correction edits tend to change the original semantics, especially for LLMs Coyne et al. (2023). To achieve this, CLEME2.0 distinguishes between necessary and unnecessary corrections and disentangles edits into four main types: true positive (TP), necessary false positive (FP), unnecessary false positive (FP), false negative (FN) edits.222True negative edits are not considered in our method. For example in Figure 1, the Hyp.1 makes three necessary edits on the right positions, where [the ] is a TP edit but two of others ([were was] and [for in]) are FP edits since they are not covered in the reference. So Hyp.1 tends to mistakenly correct grammatically errors. On the other hand, Hyp.2 makes extra two FP edits ([ of] and ([century centuries]) since the reference does not correct the right positions, indicating the occurrence of two under-correction phenomena. Additionally, [for for] of the Hyp.2 is considered as an FN edit, which means the occurrence of an under-correction phenomenon. Since the edit disentanglement is based on the chunk partition technique proposed by CLEME, so we dub this strategy as CLEME2.0.
Disentangling edits enables us to investigate concrete dimensions of GEC systems by computing upon an evaluation dataset four disentangled scores: hit-correction, error-correction, under-correction, and over-correction scores. In contrast to mainstream GEC metrics like ERRANT Bryant et al. (2017) and MaxMatch Dahlmeier and Ng (2012a) that reveal the system performance by P/R/F0.5, this disentanglement can provide an interpretable insight into fine-grained dimensions responsible for describing critical characteristics of GEC systems. Then, we integrate these disentangled scores into a comprehensive score using linear weighted summation, placing different emphases on disentangled scores. We leverage the comprehensive score to indicate the system performance from a global perspective. Similar to CLEME Ye et al. (2023c), CLEME2.0 also supports the evaluation based on either correction dependence or correction independence assumptions, providing a flexible option.
Besides, we assume that edits with various extents of modification should affect distinctively the evaluation results. Therefore, we incorporate two edit weighting techniques into CLEME2.0, namely similarity-based weighting Gong et al. (2022) and LLM-based weighting. Specifically, the techniques compute an important weight for each edit using a language model rather than treating each edit equally, thus equipping CLEME2.0 with abilities to capture context semantics and overcome the defect of traditional measures relying on superficial form similarity Kobayashi et al. (2024a).
To verify the effectiveness of CLEME2.0, we conduct extensive experiments on 2 human judgment datasets (GJG15 Grundkiewicz et al. (2015) and SEEDA Kobayashi et al. (2024b)), where our method consistently achieves high correlations. We also demonstrate the robustness of CLEME2.0 by computing the evaluation results based on 6 reference datasets with disparate annotation styles. In summary, our contributions are three folds:
We propose CLEME2.0, a more interpretable evaluation strategy, which is beneficial to reveal crucial characteristics of GEC systems.
We boost CLEME2.0 with two edit weighting techniques, including similarity-based and LLM-based weighting, to overcome the inability of traditional reference-based metrics.
Extensive experiments and analyses are conducted to confirm the effectiveness and robustness of our proposed method.
2 Related Work
Reference-based metrics.
Reference-based metrics evaluate GEC systems by referencing manually written materials Ye et al. (2023a, b); Huang et al. (2023); Li et al. (2024a). The M2 scorer Dahlmeier and Ng (2012b) identifies optimal edit sequences between source sentences and system hypotheses, using the F0.5 score. However, this method can inflate scores by manipulating edit boundaries. Bryant et al. (2017) proposed ERRANT, which improves edit extraction with a linguistically-informed alignment algorithm, but it remains language-dependent and biased in multi-reference evaluation. Napoles et al. (2015) introduced GLEU, an n-gram-based metric inspired by BLEU for GEC evaluation. Ye et al. (2023c) proposed CLEME to eliminate bias in multi-reference evaluation by transforming the source, hypothesis, and references into chunk sequences with consistent boundaries, providing unbiased F0.5 scores. Gong et al. (2022) introduce PT-M2, focusing on scoring changed words extracted by the M2 metric.
Reference-less metrics.
To overcome the limitations of reference-based metrics, recent research focus on reference-less scoring. Inspired by quality estimation in NMT Liu et al. (2022); Dong et al. (2023), Napoles et al. (2016a) propose Grammaticality-Based Metrics (GBMs) using an existing GEC system or a pre-trained ridge regression model. Asano et al. (2017) enhance GBMs by adding criteria like grammaticality, fluency, and meaning preservation. Yoshimura et al. (2020b) introduce SOME, which uses sub-metrics optimized for manual assessment with regression models. Scribendi Score Islam and Magnani (2021) combines language perplexity and token/Levenshtein distance ratios. IMPARA Maeda et al. (2022) incorporates a Quality Estimator and a Semantic Estimator based on BERT to evaluate GEC output quality and semantic similarity. While reference-less metrics align well with human judgments, they lack interpretability due to the heavy dependence on trained models, thus posing latent risks.
3 Method
Our CLEME2.0 can be generally divided into three main steps, with the overview shown in Figure 2. Additionally, we incorporate two distinct edit weighting techniques to enhance performance.
3.1 Edit Extraction
The first step is edit extraction. Given a source sentence and a target (either hypothesis or reference) sentence , this step is to extract the edits describing the modification from to . Here, we utilize the chunk partition technique from CLEME Ye et al. (2023c) to execute the process of edit extraction. Unlike the traditional metrics like ERRANT Bryant et al. (2017) and MaxMatch Dahlmeier and Ng (2012a), CLEME concurrently aligns all sentences, including the source, the hypothesis, and all the references. This facilitates segmentation of them all into chunk sequences with an equal number of chunks, irrespective of the varying token counts in different sentences, as delineated in Figure 2. It is worth noting that a chunk is a basic edit unit, which can be unchanged, corrected, or dummy (empty) Ye et al. (2023c).
3.2 Disentangled Scores
For the purpose of computing disentangled scores, we initially disentangle edits into four core types. 1) TP edits refer to the corrected/dummy hypothesis chunks that share the same tokens as the corresponding reference chunks. 2) FP edits are the corrected/dummy hypothesis chunks that have different tokens from those in the corresponding reference chunks wherein the reference chunks are also corrected/dummy ones. 3) FP edits are the corrected hypothesis chunks but their corresponding reference chunks remain unchanged. 4) FN edits indicate the unchanged hypothesis chunks but the corresponding reference chunks are corrected/dummy. It is highlighted that traditional metrics Dahlmeier and Ng (2012a); Bryant et al. (2017); Li et al. (2023b) do not distinguish between FP and FP, treating both as FP, thereby resulting in confusion between error-correction and over-correction. Actually, we have .
Furthermore, we can differentiate between necessary and unnecessary edits. TP, FP, and FN edits are all necessary edits, since their corresponding reference chunks are also corrected/dummy, implying the existence of grammatical errors in the related parts of . On the contrary, FP edit are unnecessary edits because the systems propose corrections not represented in references. Consequently, we can define four disentangled scores.
Hit-correction score.
This paper defines the hit-correction score as the ratio of TP edits to all necessary reference edits. Its purpose is to quantify the accuracy with which systems offer correct corrections. The formula is as follows:
Error-correction score.
Conversely, the error-correction score is defined as the ratio of FP edits to all necessary reference edits. This score seeks to evaluate the degree to which systems generate erroneous corrections for grammatical errors. The formula for this score is as follows:
Under-correction score.
Similarly, the under-correction score is proposed to measure the degree to which systems omit to correct grammatical errors, which is computed as follow:
Over-correction score.
The score is introduced in response to frequent observations that LLMs are prone to over-correcting texts. This score is determined by the proportion of FP edits to all hypothesis corrected/dummy edits, aiming to gauge the level to which systems offer excessive corrections:
3.3 Comprehensive Score
Once the four disentangled scores have been computed, they need to be merged into a comprehensive score that encapsulates the global performance of the systems. We employ a weighted summation approach to organize these four scores for interpretability and simplification. By definition, systems with higher hit-correction scores are usually preferable, a tendency that inversely applies to the remaining scores. Thus, the comprehensive score can be calculated using the following formula:
where is the trade-off factor for each disentangled score, and we constrain that and .
3.4 Edit Weighting
Existing reference-based metrics, such as ERRANT and CLEME, depend heavily on superficial literal similarity. This means that, regardless of length or modification, all types of edits have equal weighting in the evaluation scores. This aspect fails to acknowledge that human evaluators might semantically consider the edits’ varying importance levels. Therefore, we introduce two distinct edit weighting techniques to compute the importance weights of edits. These weights are then incorporated into the calculation of the aforementioned disentangled scores as depicted in Equation (1) (4). Take the hit-correction score as a typical example, we reformulate the Equation (1) as follow:
Similarity-based weighting.
We use PTScore from Gong et al. (2022) to provide edit weights. Through simulating a partially accurate version of the source sentence , PTScore can assign individual weights to edits, in spite of multiple edits in a sentence. Since it performs based on BERTScore Zhang et al. (2019) designed to compute similarity scores for text generation, we call this technique as similarity-based weighting. The computation process is as follows:
where the function is intended to replace a specific chunk of the source with the corrected/dummy hypothesis chunk . Here, denotes the reference sentence. Comprehensive details can be found in Gong et al. (2022).
LLM-based weighting.
In light of the impressive semantic understanding capabilities of LLMs Tan et al. (2023); Li et al. (2024b); Yu et al. (2024); Huang et al. (2024); Du et al. (2024); Xu et al. (2024); Qin et al. (2024b), some recent work has sought their use in evaluating assorted NLP tasks Pavlovic and Poesio (2024); Chen et al. (2024); Qin et al. (2024a), GEC evaluations included Sottana et al. (2023). Drawing inspiration from this trend, we employ Llama-2-7B Touvron et al. (2023) as a scorer to acquire the importance score for each edit. These scores range from 1 to 5, with higher scores indicating increased edit importance. We provide the implementation details and the prompting template in Appendix D.
4 Experiments
4.1 Experimental Settings
| Metric | CoNLL-2014 | BN-10GEC | E-Minimal | E-Fluency | NE-Minimal | NE-Fluency | Avg. | |||||||
| EW | TS | EW | TS | EW | TS | EW | TS | EW | TS | EW | TS | |||
| M2 | 0.623 | 0.672 | 0.547 | 0.610 | 0.597 | 0.650 | 0.590 | 0.659 | 0.575 | 0.634 | 0.582 | 0.649 | 0.616 | |
| 0.687 | 0.720 | 0.648 | 0.692 | 0.654 | 0.703 | 0.654 | 0.709 | 0.577 | 0.648 | 0.648 | 0.703 | 0.670 | ||
[…]
| 0.467 | 0.555 | 0.754 | 0.806 | 0.577 | 0.511 | 0.710 | 0.757 | -0.005 | -0.055 | 0.725 | 0.819 | 0.551 | ||
[…]
| 0.659 | 0.698 | 0.637 | 0.698 | 0.742 | 0.786 | 0.720 | 0.775 | 0.747 | 0.797 | 0.753 | 0.797 | 0.734 | ||
[…]
| 0.758 | 0.769 | 0.690 | 0.824 | 0.709 | 0.736 | 0.758 | 0.802 | 0.736 | 0.758 | 0.758 | 0.802 | 0.758 | ||
[…]
| 0.709 | 0.742 | 0.692 | 0.747 | 0.797 | 0.813 | 0.714 | 0.775 | 0.786 | 0.835 | 0.720 | 0.791 | 0.760 | ||
[…]
| 0.709 | 0.731 | 0.692 | 0.747 | 0.791 | 0.802 | 0.731 | 0.791 | 0.797 | 0.841 | 0.714 | 0.786 | 0.761 | ||
[…]
| 0.665 | 0.736 | 0.626 | 0.692 | 0.736 | 0.808 | 0.742 | 0.830 | 0.775 | 0.846 | 0.599 | 0.714 | 0.730 | ||
[…]
| 0.665 | 0.736 | 0.698 | 0.758 | 0.736 | 0.808 | 0.742 | 0.830 | 0.775 | 0.846 | 0.670 | 0.769 | 0.753 | ||
[…]
| 0.819 | 0.890 | 0.802 | 0.863 | 0.791 | 0.868 | 0.758 | 0.852 | 0.830 | 0.896 | 0.786 | 0.857 | 0.834 | ||
[…]
| 0.846 | 0.901 | 0.835 | 0.885 | 0.819 | 0.885 | 0.758 | 0.852 | 0.846 | 0.896 | 0.863 | 0.923 | 0.859 | ||
| SentM2 | 0.871 | 0.864 | 0.567 | 0.646 | 0.805♣ | 0.836♣ | 0.655 | 0.732 | 0.729♣ | 0.785♣ | 0.621 | 0.699 | 0.734 | |
| 0.731 | 0.758 | 0.593 | 0.648 | 0.806♣ | 0.845♣ | 0.731 | 0.764 | 0.797♣ | 0.846♣ | 0.632 | 0.687 | 0.737 | ||
[…]
| 0.720 | 0.775 | 0.769 | 0.824 | 0.764♣ | 0.797♣ | 0.791 | 0.846 | 0.764♣ | 0.830♣ | 0.768 | 0.846 | 0.791 | ||
[…]
| 0.742 | 0.747 | 0.786 | 0.830 | 0.775♣ | 0.819♣ | 0.813 | 0.764 | 0.780♣ | 0.841♣ | 0.830 | 0.857 | 0.799 | ||
[…]
| 0.907 | 0.874 | 0.626♣ | 0.670♣ | 0.808♣ | 0.819♣ | 0.797 | 0.841 | 0.813♣ | 0.857♣ | 0.742 | 0.786 | 0.795 | ||
[…]
| 0.824 | 0.808 | 0.835 | 0.874 | 0.775♣ | 0.819♣ | 0.824 | 0.863 | 0.797♣ | 0.846♣ | 0.791 | 0.846 | 0.825 | ||
[…]
| 0.725 | 0.758 | 0.659♣ | 0.714♣ | 0.775♣ | 0.819♣ | 0.808 | 0.846 | 0.819♣ | 0.874♣ | 0.762 | 0.825 | 0.782 | ||
[…]
| 0.714 | 0.725 | 0.681 | 0.747 | 0.857♣ | 0.885♣ | 0.824 | 0.901 | 0.857♣ | 0.912♣ | 0.720 | 0.791 | 0.801 | ||
[…]
| 0.709 | 0.720 | 0.681 | 0.747 | 0.879♣ | 0.912♣ | 0.857 | 0.923 | 0.824♣ | 0.885♣ | 0.654 | 0.720 | 0.793 | ||
[…]
| 0.907 | 0.912 | 0.808 | 0.863 | 0.852♣ | 0.879♣ | 0.885 | 0.945 | 0.753♣ | 0.780♣ | 0.896 | 0.940 | 0.868 | ||
[…]
| 0.868 | 0.879 | 0.753 | 0.824 | 0.863♣ | 0.901♣ | 0.879 | 0.956 | 0.775♣ | 0.802♣ | 0.835 | 0.923 | 0.855 | ||
Human ranking datasets.
We conduct comprehensive experiments across two human judgment datasets with disparate annotation protocols.
SEEDA. Kobayashi et al. (2024b) reveal several shortcomings in GJS15 and subsequently propose SEEDA, an alternative dataset featuring human judgments across two levels of granularity. To align with the contemporary trend in GEC, SEEDA is primarily focused on mainstream neural-based systems.
Both of human judgment datasets derive the overall human rankings for all GEC systems by employing Expected Wins (EW) Bojar et al. (2013) and TrueSkill (TS) Sakaguchi et al. (2014) methods. Following the previous approaches Ye et al. (2023c); Kobayashi et al. (2024b), we compute the Pearson () and Spearman () correlations between metrics and human judgments, in order to ascertain the effectiveness and robustness of GEC metrics within the context of system-level ranking.
Reference datasets.
Reference-based metrics rely on a reference set to establish a system ranking list, the properties of which may significantly influence the performance of the metrics. To investigate the impact of variable reference sets, we assess human consistency across 6 reference datasets. These datasets encompass a range of annotation styles, and a number of human annotators, including CoNLL-2014 Grundkiewicz et al. (2015), BN-10GEC Bryant and Ng (2015) and SN-8GEC Sakaguchi et al. (2016). Notably, SN-8GEC is partitioned into 4 sub-sets, namely E-Minimal, E-Fluency, NE-Minimal, and NE-Fluency. A more thorough breakdown of these datasets and the statistics are provided in Appendix A.
Corpus and sentence levels.
GEC evaluation metrics can compute an overall system-level score for a given system in two settings Gong et al. (2022). Given the metric , source sentences , hypothesis sentences and reference sentences , 1) corpus-level metrics compute the system score based on the whole corpus , and 2) sentence-level metrics use the average of the sentence-level scores .
| Metric | SEEDA-S | SEEDA-E | Avg. | ||
| M2 | 0.658 | 0.487 | 0.791 | 0.764 | 0.675 |
| PT-M2 | 0.845 | 0.769 | 0.896 | 0.909 | 0.855 |
| ERRANT | 0.557 | 0.406 | 0.697 | 0.671 | 0.583 |
| PT-ERRANT | 0.818 | 0.720 | 0.888 | 0.888 | 0.829 |
| GoToScorer | 0.929 | 0.881 | 0.901 | 0.937 | 0.912 |
| GLEU | 0.847 | 0.886 | 0.911 | 0.897 | 0.885 |
| Scribendi Score | 0.631 | 0.641 | 0.830 | 0.848 | 0.738 |
| SOME | 0.892 | 0.867 | 0.901 | 0.951 | 0.903 |
| IMPARA | 0.911 | 0.874 | 0.889 | 0.944 | 0.903 |
| CLEME-dep | 0.633 | 0.501 | 0.755 | 0.757 | 0.662 |
| CLEME-ind | 0.616 | 0.466 | 0.736 | 0.708 | 0.632 |
| CLEME2.0-dep (Ours) | 0.937 | 0.865 | 0.945 | 0.939 | 0.922 |
| CLEME2.0-ind (Ours) | 0.908 | 0.844 | 0.961 | 0.946 | 0.915 |
| CLEME2.0-sim-dep (Ours) | 0.923 | 0.914 | 0.948 | 0.974 | 0.940 |
| CLEME2.0-sim-ind (Ours) | 0.921 | 0.907 | 0.953 | 0.981 | 0.941 |
| Sent-M2 | 0.802 | 0.692 | 0.887 | 0.846 | 0.807 |
| SentERRANT | 0.758 | 0.643 | 0.860 | 0.825 | 0.772 |
| SentCLEME-dep | 0.866 | 0.809 | 0.944 | 0.939 | 0.890 |
| SentCLEME-ind | 0.864 | 0.858 | 0.935 | 0.911 | 0.892 |
| SentCLEME2.0-dep (Ours) | 0.905 | 0.844 | 0.955 | 0.946 | 0.913 |
| SentCLEME2.0-ind (Ours) | 0.875 | 0.837 | 0.953 | 0.953 | 0.905 |
| SentCLEME2.0-sim-dep (Ours) | 0.924 | 0.858 | 0.923 | 0.953 | 0.915 |
| SentCLEME2.0-sim-ind (Ours) | 0.921 | 0.886 | 0.957 | 0.960 | 0.931 |
| Metric | EW | TS | Avg. | ||
| CLEME2.0-dep-Hit | 0.599 | 0.593 | 0.673 | 0.648 | 0.628 |
| CLEME2.0-dep-Error | -0.444 | -0.533 | -0.526 | -0.593 | -0.524 |
| CLEME2.0-dep-Under | 0.496 | 0.599 | 0.576 | 0.659 | 0.583 |
| CLEME2.0-dep-Over | 0.118 | 0.269 | 0.073 | 0.275 | 0.253 |
| SentCLEME2.0-dep-Hit | 0.594 | 0.593 | 0.672 | 0.648 | 0.627 |
| SentCLEME2.0-dep-Error | -0.405 | -0.429 | -0.489 | -0.500 | -0.456 |
| SentCLEME2.0-dep-Under | 0.489 | 0.511 | 0.572 | 0.582 | 0.539 |
| SentCLEME2.0-dep-Over | -0.247 | -0.363 | -0.346 | -0.440 | -0.349 |
Trade-off factors.
We employ a cross-evaluation approach to determine two optimal configurations for trade-off factors applicable at the corpus and sentence levels, respectively. At the corpus level, we assign the factors as . Conversely, at the sentence level, these are adjusted to . Additionally, CLEME2.0 metrics named with “sim” are based on similarity-based edit weighting, and we leave the analysis of LLM-based edit weighting to Section 5.2.
Evaluation Assumptions.
Ye et al. (2023c) propose evaluating systems based on one of two assumptions, namely correction dependence and correction independence. In short, the correction independence assumption offers a more relaxed edit matching process, implying that systems are more inclined to yield higher scores when multiple references are accessible. Inspired by this work, CLEME2.0 also supports both assumptions, and we will study their impact on our method.
4.2 Results of GJG15 Ranking
The correlations between the GEC metrics and human judgments on the GJG15 rankings are shown in Table 1, and we have the following insights.
CLEME2.0 outperforms other metrics at both corpus and sentence levels.
For corpus-level, CLEME2.0-sim-ind achieves the highest average correlations, closely followed by CLEME2.0-sim-dep. CLEME2.0-ind and CLEME2.0-dep can also gain comparable correlations with other metrics, in spite of the fact that they do not utilize any edit weighting techniques. On the other hand, sentence-level metrics exhibit a similar pattern. SentCLEME2.0-sim-dep and SentCLEME2.0-sim-ind achieve the highest Pearson and Spearson correlations, respectively. These results significantly demonstrate the effectiveness and robustness of our proposed method across different settings.
Sentence-level metrics outperform their corpus-level counterparts.
This observation aligns with recent studies Gong et al. (2022); Ye et al. (2023c). This is because system-level rankings treat each sample equally, which is consistent with the evaluation process of sentence-level metrics. In contrast, corpus-level metrics allow samples with more edits to significantly influence the results. SentPT-M2 exhibits the best performance on CoNLL-2014, but its results on BN-10GEC, E-Minimal, and NE-Fluency are inferior compared to our method.
In general, our method aligns more consistently with human judgments than existing mainstream metrics. Particularly, most weighted outcomes outshine the unweighted ones, attributable to the incorporation of semantic considerations. However, on E-Minimal and NE-Minimal, the unweighted and weighted results exhibit comparability. We conjecture that this could be due to the annotations in these datasets being minimal yet decisive, reducing the possibility of generating diverse weights.
4.3 Results of SEEDA Ranking
We conduct a supplementary experiment on the SEEDA-Sentence and SEEDA-Edit datasets, comparing our method with a wide range of GEC metrics. Table 2 demonstrates again that our approach obtains the optimal results on both datasets. Kobayashi et al. (2024b) claim that the correlations of most metrics tend to decline when shifting from classical to neural systems in evaluation. This suggests that traditional metrics may struggle when assessing more extensively edited and fluent corrections generated by neural systems. However, our method is still able to address these challenges and obtain even better results. The results on SEEDA-Edit surpass those on SEEDA-Sentence due to the finer granularity of SEED-Edit, which is more consistent with the functioning of CLEME2.0.
It is crucial to mention that reference-less metrics such as SOME and IMPARA yield high outcomes, in part, because these are fine-tuned on GEC data. Although fine-tuned metrics generally perform better, they are not without their limitations. Firstly, the incorporation of fine-tuning in SOME and IMPARA makes these reference-less metrics more costly. Second, these reference-less metrics may suffer from poor robustness since the assessment process is not guided by human-annotated references. For example, the authors of Scribendi Score claim that it can achieve high correlations on the human judgment dataset from Napoles et al. (2016b). However, only moderate correlations are observable on SEEDA-Edit.
| Metric | EW | TS | Avg. | ||
| CLEME2.0-dep | 0.461 | 0.423 | 0.483 | 0.457 | 0.456 |
| CLEME2.0-ind | 0.468 | 0.421 | 0.489 | 0.453 | 0.458 |
| CLEME2.0-sim-dep | 0.559 | 0.592 | 0.581 | 0.624 | 0.589 |
| CLEME2.0-sim-ind | 0.566 | 0.593 | 0.588 | 0.622 | 0.592 |
| SentCLEME2.0-dep | 0.374 | 0.305 | 0.362 | 0.290 | 0.333 |
| SentCLEME2.0-ind | 0.372 | 0.302 | 0.356 | 0.283 | 0.328 |
| SentCLEME2.0-sim-dep | 0.410 | 0.361 | 0.400 | 0.345 | 0.379 |
| SentCLEME2.0-sim-ind | 0.412 | 0.360 | 0.399 | 0.338 | 0.377 |
| Dataset | Corpus-EW | Corpus-TS | Sentence-EW | Sentence-TS | ||||
| CoNLL-2014 | 0.697 | 0.659 | 0.759 | 0.720 | 0.626 | 0.654 | 0.696 | 0.698 |
| BN-10GEC | 0.732 | 0.764 | 0.796 | 0.813 | 0.638 | 0.637 | 0.708 | 0.698 |
| E-Minimal | 0.709 | 0.786 | 0.779 | 0.819 | 0.642 | 0.692 | 0.715 | 0.747 |
| E-Fluency | 0.760 | 0.786 | 0.831 | 0.841 | 0.642 | 0.665 | 0.720 | 0.714 |
| NE-Minimal | 0.777 | 0.823 | 0.839 | 0.861 | 0.654 | 0.747 | 0.723 | 0.791 |
| NE-Fluency | 0.823 | 0.692 | 0.849 | 0.709 | 0.664 | 0.791 | 0.742 | 0.830 |
| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | Chunk 7 | Chunk 8 | Chunk 9 | |
| Source | When we are | diagonosed out | with certain genetic | disease | , are we suppose to disclose | this result | to | our | relatives ? |
| Ref. | When we are | diagnosed | with certain genetic | diseases | , are we suppose to disclose | this result | to | our | relatives ? |
| Hyp. | When we are | diagnosed out (0.056) | with certain genetic | diseases (0.006) | , are we suppose to disclose | the results (0.019) | to | their (0.021) | relatives ? |
| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | |
| Source | Do one | who | suffered | from this disease keep it a secret | of infrom | their relatives ? |
| Ref | Does one | who | suffers | from this disease keep it a secret | or inform | their relatives ? |
| Hyp. | Do one (0.028) | who | suffer (0.011) | from this disease keep it a secret | to inform (0.094) | their relatives ? |
5 Analysis
5.1 Ablation Studies
Isolated effect of each disentangled score.
We perform ablation experiments on (Sent)CLEME2.0-dep to observe how each disentangled score performs. Since a system exhibiting lower error-correction, under-correction, and over-correction is more desirable, these scores are subtracted from 1. The results are presented in Table 3. Both hit-correction and under-correction scores display moderate correlations. Over-correction scores exhibit slight positive correlations at the corpus-level, but negligible negative correlations at the sentence level. Interestingly, error-correction scores manifest negative correlations. However, this does not imply that error-correction scores have no influence on the comprehensive score. In fact, the trade-off factor of error-correction scores is relatively significant. It is hypothesized that evaluations based solely on error-correction scores could unduly encourage systems that produce only highly confident edits, resulting in potential evaluation bias.
Average correlations.
To further compare different metrics from a global perspective, we report the average correlations obtained through the exhaustive enumeration of various parameter configurations. Specifically, all possible parameter combinations are attempted, with a step increment of 0.05. From Table 4, we observe that all the correlations are positive, regardless of the applied correction assumptions, evaluation levels, and weighting techniques. Comparing the unweighted and similarity-based weighted results, we conclude that similarity-based weighting significantly promotes human correlations, even on a global scale. Furthermore, corpus-level metrics tend to attain higher average results than sentence-level metrics; nonetheless, sentence-level metrics with optimal parameters surpass their corpus-level counterparts. This suggests that corpus-level metrics may exhibit enhanced robustness concerning parameter selection.
5.2 Exploration of LLM
The results of LLM-based edit weighting are shown in Table 5. The corpus-level results are quite satisfying and are comparable to those of most metrics such as PT-M2 and CLEME. However, the sentence-level outcomes are less satisfactory. This could be due to the fact that LLM assigns error-editing scores on a scale of 1 to 5, which is notably coarser when contrasted with the 0 to 1 continuous scoring scale. Sentence-level scores depend on the mean of the editing scores within a particular sentence. Consequently, even the slightest bias in the scores assigned by the LLM might lead to significant deviations in the sentence-level outcomes.
Although the LLM has had some success, its performance still falls short when compared to the similarity-based weighting. This might be due to the scoring granularity of the 1 to 5 scale provided by the LLM not being sufficiently fine-tuned. In addition, the score heavily relies on the functionality of the LLM, which proves rather unstable.
5.3 Case Study
Table 6 presents examples of CLEME2.0. In the top group, chunk 2 achieves the highest score (0.056), highlighting its significant impact on the sentence. Although "diagnosed" was correctly amended, the omission of "out" persists, rendering the sentence still incorrect. Chunk 4 represents a hit-correction relating to the singular and plural forms in the source sentence and its low score indicates a minimal impact. Chunks 6 and 8 are types of over-correction. Chunk 6 does not change the original meaning, whereas chunk 8 introduces a more severe error due to an incorrect personal pronoun. In the second group, both chunks 3 and 5 exhibit error-corrections, with chunk 5 scoring higher than chunk 3. Chunk 3 involves issues of tense and singular-plural, while chunk 5 presents a more serious error that completely alters the meaning of the sentence. The weighting scores reflect the superiority of the method. For metrics that do not apply weightings, sorts of edits are uniformly assigned, which does not reflect the actual semantics.
6 Conclusion
This paper proposes CLEME2.0, an interpretable evaluation strategy, which are beneficial to reveal crucial characteristics of GEC systems. To address the limitations of traditional reference-based metrics in capturing semantic nuances, we enhance CLEME2.0 using two innovative edit weighting techniques: similarity-based and LLM-based weighting. Through comprehensive experiments and analyses, we validate the efficacy and reliability of our approach. We believe CLEME2.0 will provide a promising perspective on the task of grammatical error correction.
Limitation
Although CLEME2.0 can be extended to other languages, its effectiveness has not been tested in any language other than English. Furthermore, all reference sets utilized in our experiments are derived from the CoNLL-2014 shared task, which is a second language dataset. To demonstrate the robustness of our approaches, further experiments on evaluation datasets encompassing multiple languages and text domains are required. Finally, we strongly advocate for the construction of new GEC evaluation datasets to advance the development of NLP.
Ethics Statement
In this paper, we validate the effectiveness and robustness of our proposed approach using the CoNLL-2014, BN-10GEC, and SN-8GEC reference datasets. These datasets are sourced from publicly available resources on legitimate websites and do not contain any sensitive data. Additionally, all the baselines employed in our experiments are publicly accessible GEC metrics, and we have duly cited the respective authors. We confirm that all datasets and baselines utilized in our experiments are consistent with their intended purposes.
Appendix A Details about GEC Meta-Evaluation
A.1 Human Rankings
GJG15 ranking.
Grundkiewicz et al. (2015) propose the first large-scale human judgement dataset towards 12 participating systems of the CoNLL-2014 shared task. In this assessment, 8 native speaker are asked to rank the outputs of all the systems from best to worst. Two system ranking lists are generated using Expected Wins (EW) and TrueSkill (TS) respectively. Since all the involved systems are mostly classical systems such as statistical machine translation approaches Junczys-Dowmunt and Grundkiewicz (2014) and classifier-based approaches Rozovskaya et al. (2014), the dataset may not be a ideal test bed for meta-evaluation in the current time.
SEEDA ranking.
Kobayashi et al. (2024b) identify several limitations of GJG15 ranking dataset, and propose a new human ranking dataset called SEEDA. SEEDA consists of corrections with human ratings along two different granularities: edit-based and sentence-based, covering 12 state-of-the-art systems including large language models (LLMs), and two human corrections with different focuses. Three native English speakers participate in the annotation process. Similar to Grundkiewicz et al. (2015), the overall human rankings are derived from TrueSkill (TS) and Expected Wins (EW) based on pairwise judgments.
| Item | CoNLL-2014 | BN-10GEC | E-Minimal | E-Fluency | NE-Minimal | NE-Fluency |
| # Sents (Length) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) |
| # Refs (Length) | 2,624 (22.8) | 13,120 (22.9) | 2,624 (23.2) | 2,624 (22.8) | 2,624 (23.0) | 2,624 (22.2) |
| # Edits (Length) | 5,937 (1.0) | 36,677 (1.0) | 4,500 (1.0) | 8,373 (1.1) | 4,964 (0.9) | 11,033 (1.2) |
| # Unchanged Chunks (Length) | 11,174 (4.8) | 93,496 (2.5) | 8,887 (6.3) | 12,823 (3.8) | 10,748 (5.1) | 14,086 (2.9) |
| # Corrected/Dummy Chunks (Length) | 4,994 (1.3) | 26,948 (2.4) | 3,963 (1.2) | 6,305 (1.7) | 4,221 (1.2) | 6,892 (2.6) |
Appendix B Statistics of Reference Datasets
Table 7 presents the statistics of all the reference sets involved in our experiments.
B.1 Baseline Metrics
In our evaluation, we compare our method with the following reference-based baseline metrics, including corpus and sentence-level variants:
M2 and SentM2 Dahlmeier and Ng (2012a) dynamically extract the hypothesis edits with the maximum overlap of gold annotations by utilizing the Levenshtein algorithm.
GLEU and SentGLEU Napoles et al. (2015) are BLEU-like GEC metrics based on n-gram matching, rewarding hypothesis n-grams that align with the reference but not the source, while penalizing those aligning solely with the source. GLEU is the main metric in JFLEG, an English GEC dataset that highlights holistic fluency edits.
ERRANT and SentERRANT Bryant et al. (2017) are the most mainstream GEC metrics, which are based on edit matching. They are able to extract edits more accurately, by utilizing the linguistically enhanced Damerau-Levenshtein algorithm.
PT-M2 and SentPT-M2 Gong et al. (2022) leverage pre-trained language model (PLM) to evaluate GEC systems. The main idea is similar to M2 and ERRANT, but they can leverage the knowledge of pre-trained language models to score edits effectively.
CLEME and SentCLEME Ye et al. (2023c) are proposed to provide unbiased scores for multi-reference evaluation. Besides, the authors introduce the correction independence assumption, so CLEME can perform based on either traditional correction dependence or correction independence assumptions.
In addition, for the evaluation of SEEDA, we have additionally added the following evaluation methods in accordance with the evaluation methods reported in Kobayashi et al. (2024b):
GoToScorer Gotou et al. (2020): takes into account the difficulty of error correction when calculating the evaluation score. The difficulty is calculated based on the number of systems that can correct errors.
SOME Yoshimura et al. (2020b): optimizes human evaluation by fine-tuning BERT separately for criteria such as grammaticality, fluency, and meaning preservation.
IMPARA Maeda et al. (2022): incorporates a quality assessment model fine-tuned using BERT parallel data and a similarity model that takes into account the effects of editing.
Appendix C Detailed Results of Evaluation
We list detailed evaluation results of CLEME2.0 on CoNLL-2014 in Table 8.
Appendix D Experimental Details of LLM-based Edit Weighting
Prompt:
As a grammar correction evaluator, you are required to score the corrected editors for each grammatical error. We will give three lines, the first line is the original sentence given, the second line is the modification made to the editor, and the third line is the output form.
The scoring range is 1-5. The larger the score, the more important the editor’s correction is. Correspondingly, the smaller the score, the less important the editor’s correction is.
1 point indicates that this editor’s modification has almost no impact on the original sentence and is dispensable.
2 points indicates that this editorial change has a slight impact.
3 points indicates that this editor’s changes have a certain impact.
4 points indicates that this editorial change is necessary.
5 points indicates that this editing modification is very necessary and of high importance.
The output format is a score of 1 to 5 points.
Next, I will give you a sentence only with an edit. You need to rate each edit in sequence. The desired output is just a score, without any redundant explanation.
Example Input:
Sentence: Nowadays the technologies were improved a lot compared to the last century.
Edit: were => have
Output (1-5):
Example Output:
5
Note that the output must be a number between 1 and 5. Here is the sample:
Due to the strong semantic understanding capabilities of large language models (LLMs) Yang et al. (2020, 2022b, 2022a); Chai et al. (2024), recent work Sottana et al. (2023) has sparked interest in using LLMs for text evaluation, including the evaluation of grammatical error correction. Inspired by this, we utilize LLMs as weighted scorers to assess the importance of each edit. The template for the LLM is shown in Figure 3. For each edit, the constructed sentence contains only one grammatical error, while the other positions are correct. The second line shows the modification of that edit. The LLM is required to determine the necessity of the modified edit and output a score from 1 to 5. A higher score indicates a greater necessity for the edit modification. We do not inform the LLM of the specific types of edits; instead, we let the larger model evaluate the necessity of the modified edits.
D.1 Hit-Correction Edits
Scenario: The hypothesis and reference sentence are consistent.
Focus: The significance of the transition from the source to the hypothesis sentence.
Scoring: A higher score indicates that the edit from source to reference sentence carries substantial importance. Conversely, a lower score suggests that this transition is less crucial.
D.2 Error-Correction Edits
Scenario: The hypothesis and reference sentence are inconsistent.
Focus: The significance of the transition from the hypothesis to the reference sentence.
Scoring: A high score indicates a critical edit, suggesting significant inaccuracies in the hypothesis sentence. A low score implies that the modification is of minimal importance, indicating the hypothesis sentence is either correct or not substantially incorrect.
D.3 Under-Correction Edits
Scenario: The source and hypothesis sentence remain unchanged.
Focus: The importance of modifications from the source to the reference sentence.
Scoring: A high score implies a critical need for the edit, pointing to a severe under-correction. Conversely, a low score indicates that the edit is of lesser importance, suggesting a mild under-correction.
D.4 Over-Correction Edits
Scenario: The source is equivalent to the reference sentence, leading to two distinct situations:
1.
The reference is not an ideal sentence, and the hypothesis sentence is corrected but deemed overcorrected.
2.
The reference is optimal, necessitating no amendments, yet the hypothesis sentence introduces corrections.
Evaluation:
First Situation: Assess the importance () of the edit from the source to the hypothesis sentence. A higher W1 score indicates that the edit is crucial, suggesting imperfections in the reference sentence. Conversely, a lower score suggests that the edit is of minimal importance, rendering the hypothesis’s correction unnecessary.
Second Situation: Examine the significance () of the edit from the hypothesis to the reference sentence. A higher score indicates that the edit is critical, denoting that the hypothesis’s correction was overly aggressive. A lower score implies the edit was unneeded, making the correction by the hypothesis irrelevant.
Formula: The computation of over-correction score is defined as follow:
This score can be either positive or negative. A higher over-correction score signals a less effective performance by the correction system.
By systematically assessing the necessity and importance of different types of edits, we can better understand and improve the performance of grammatical error correction systems.
| Metric | AMU | CAMB | CUUI | IITB | INPUT | IPN | NTHU | PKU | POST | RAC | SJTU | UFC | UMC | |
| CLEME2.0-dependent | TP | 380 | 584 | 471 | 22 | 0 | 39 | 330 | 246 | 412 | 254 | 85 | 32 | 260 |
| sim | 9.20 | 12.66 | 7.58 | 0.39 | 0.00 | 0.77 | 5.79 | 6.69 | 8.80 | 6.68 | 1.50 | 0.42 | 5.29 | |
| FP | 817 | 1307 | 964 | 67 | 0 | 488 | 905 | 709 | 1145 | 782 | 272 | 18 | 789 | |
| sim | 16.03 | 30.92 | 16.06 | 1.80 | 0.00 | 11.93 | 24.56 | 14.36 | 19.25 | 11.98 | 6.49 | 0.25 | 18.26 | |
| FPne | 276 | 418 | 311 | 34 | 0 | 149 | 302 | 254 | 316 | 259 | 76 | 12 | 245 | |
| sim | 4.08 | 6.55 | 3.68 | 0.75 | 0.00 | 4.61 | 5.89 | 4.06 | 4.60 | 3.80 | 2.30 | 0.17 | 3.83 | |
| FPun | 541 | 889 | 653 | 33 | 0 | 339 | 603 | 455 | 829 | 523 | 196 | 6 | 544 | |
| sim | 11.95 | 24.36 | 12.38 | 1.05 | 0.00 | 7.33 | 18.67 | 10.30 | 14.64 | 8.18 | 4.19 | 0.08 | 14.43 | |
| FN | 1360 | 1150 | 1357 | 2057 | 1782 | 2886 | 1388 | 1454 | 1354 | 1487 | 1668 | 2087 | 1461 | |
| sim | 34.25 | 28.45 | 36.21 | 78.39 | 48.24 | 83.10 | 46.53 | 36.15 | 34.48 | 38.00 | 56.28 | 51.27 | 39.60 | |
| TN | 6298 | 6329 | 6224 | 6007 | 6308 | 5160 | 6274 | 6286 | 6428 | 6276 | 6313 | 5965 | 6355 | |
| sim | 6237 | 6301 | 6190 | 5373 | 6226 | 5241 | 5973 | 6214 | 6382 | 6194 | 5902 | 6092 | 6273 | |
| Hit | 0.188 | 0.271 | 0.220 | 0.010 | 0.00 | 0.013 | 0.163 | 0.126 | 0.198 | 0.127 | 0.046 | 0.015 | 0.132 | |
| sim | 0.194 | 0.266 | 0.160 | 0.005 | 0.00 | 0.009 | 0.100 | 0.143 | 0.184 | 0.138 | 0.025 | 0.008 | 0.109 | |
| Error | 0.137 | 0.194 | 0.145 | 0.016 | 0.00 | 0.048 | 0.150 | 0.130 | 0.152 | 0.130 | 0.042 | 0.006 | 0.125 | |
| sim | 0.086 | 0.138 | 0.078 | 0.009 | 0.00 | 0.052 | 0.101 | 0.0866 | 0.096 | 0.078 | 0.038 | 0.003 | 0.079 | |
| Under | 0.675 | 0.534 | 0.634 | 0.973 | 1.00 | 0.939 | 0.687 | 0.744 | 0.650 | 0.744 | 0.912 | 0.979 | 0.743 | |
| sim | 0.721 | 0.597 | 0.763 | 0.986 | 1.00 | 0.939 | 0.799 | 0.771 | 0.720 | 0.784 | 0.937 | 0.989 | 0.813 | |
| Over | 0.452 | 0.470 | 0.455 | 0.371 | 0.00 | 0.643 | 0.488 | 0.476 | 0.532 | 0.505 | 0.549 | 0.12 | 0.519 | |
| sim | 0.474 | 0.559 | 0.524 | 0.478 | 0.00 | 0.577 | 0.615 | 0.490 | 0.522 | 0.438 | 0.524 | 0.116 | 0.613 | |
| Score | 0.483 | 0.508 | 0.497 | 0.431 | 0.45 | 0.408 | 0.463 | 0.450 | 0.479 | 0.505 | 0.434 | 0.450 | 0.453 | |
| sim | 0.503 | 0.520 | 0.484 | 0.425 | 0.45 | 0.408 | 0.439 | 0.474 | 0.491 | 0.438 | 0.424 | 0.448 | 0.452 | |
[…]
| sim | 9.14 | 12.63 | 7.52 | 0.39 | 0.00 | 0.76 | 5.72 | 6.65 | 8.75 | 6.59 | 1.48 | 0.42 | 5.23 | |
| FP | 821 | 1311 | 968 | 67 | 0 | 488 | 908 | 711 | 1148 | 785 | 273 | 18 | 790 | |
| sim | 16.49 | 31.25 | 16.50 | 1.85 | 0.00 | 13.00 | 24.83 | 14.38 | 19.36 | 12.34 | 7.13 | 0.26 | 18.47 | |
| FPne | 286 | 431 | 320 | 22 | 0 | 132 | 310 | 262 | 326 | 271 | 81 | 10 | 255 | |
| sim | 4.60 | 7.51 | 4.27 | 0.44 | 0.00 | 2.62 | 6.58 | 4.58 | 5.06 | 4.02 | 1.28 | 0.15 | 4.39 | |
| FPun | 535 | 880 | 648 | 45 | 0 | 356 | 598 | 449 | 822 | 514 | 192 | 8 | 535 | |
| sim | 11.89 | 23.74 | 12.23 | 1.42 | 0.00 | 10.39 | 18.24 | 9.80 | 14.30 | 8.32 | 5.85 | 0.12 | 14.07 | |
| FN | 1600 | 1374 | 1577 | 1972 | 1982 | 1940 | 1660 | 1712 | 1587 | 1744 | 1900 | 1980 | 1714 | |
| sim | 43.65 | 35.92 | 45.22 | 57.46 | 58.31 | 54.69 | 46.92 | 46.02 | 43.09 | 46.05 | 55.32 | 58.35 | 48.02 | |
| TN | 6058 | 6105 | 6004 | 6092 | 6108 | 6106 | 6002 | 6028 | 6195 | 6019 | 6081 | 6072 | 6102 | |
| sim | 6052 | 6095 | 6009 | 6093 | 6106 | 6115 | 5995 | 6012 | 6203 | 6027 | 6079 | 6070 | 6115 | |
| Hit | 0.136 | 0.210 | 0.163 | 0.008 | 0.00 | 0.013 | 0.119 | 0.088 | 0.142 | 0.089 | 0.032 | 0.012 | 0.091 | |
| sim | 0.131 | 0.205 | 0.142 | 0.007 | 0.00 | 0.011 | 0.104 | 0.088 | 0.129 | 0.086 | 0.027 | 0.008 | 0.087 | |
| Error | 0.080 | 0.129 | 0.090 | 0.005 | 0.00 | 0.038 | 0.095 | 0.076 | 0.088 | 0.071 | 0.023 | 0.002 | 0.070 | |
| sim | 0.063 | 0.102 | 0.066 | 0.004 | 0.00 | 0.033 | 0.079 | 0.059 | 0.070 | 0.051 | 0.020 | 0.001 | 0.059 | |
| Under | 0.500 | 0.392 | 0.479 | 0.675 | 0.687 | 0.639 | 0.496 | 0.538 | 0.486 | 0.551 | 0.637 | 0.678 | 0.546 | |
| sim | 0.519 | 0.419 | 0.517 | 0.673 | 0.684 | 0.645 | 0.524 | 0.553 | 0.509 | 0.567 | 0.641 | 0.680 | 0.557 | |
| Over | 0.248 | 0.419 | 0.293 | 0.031 | 0.00 | 0.242 | 0.304 | 0.235 | 0.342 | 0.232 | 0.121 | 0.006 | 0.267 | |
| sim | 0.241 | 0.421 | 0.294 | 0.030 | 0.00 | 0.224 | 0.302 | 0.224 | 0.331 | 0.203 | 0.119 | 0.005 | 0.267 | |
| Score | 0.498 | 0.513 | 0.507 | 0.467 | 0.466 | 0.447 | 0.481 | 0.475 | 0.495 | 0.477 | 0.469 | 0.471 | 0.476 | |
| sim | 0.502 | 0.520 | 0.504 | 0.467 | 0.466 | 0.449 | 0.479 | 0.481 | 0.494 | 0.484 | 0.467 | 0.469 | 0.479 | |
| CLEME2.0-independent | TP | 388 | 596 | 487 | 22 | 0 | 39 | 338 | 248 | 420 | 255 | 85 | 32 | 262 |
| sim | 9.47 | 13.11 | 7.99 | 0.40 | 0.00 | 0.81 | 6.13 | 6.80 | 9.07 | 6.91 | 1.54 | 0.47 | 5.49 | |
| FP | 809 | 1295 | 948 | 67 | 0 | 488 | 897 | 707 | 1137 | 781 | 272 | 18 | 787 | |
| sim | 14.74 | 28.11 | 14.42 | 1.91 | 0.00 | 11.82 | 22.93 | 13.03 | 17.62 | 11.23 | 6.46 | 0.25 | 16.99 | |
| FPne | 408 | 627 | 449 | 34 | 0 | 234 | 447 | 388 | 487 | 406 | 134 | 12 | 366 | |
| sim | 6.32 | 10.62 | 5.51 | 0.86 | 0.00 | 4.79 | 9.50 | 7.30 | 7.12 | 5.56 | 2.41 | 0.17 | 6.14 | |
| FPun | 401 | 668 | 499 | 33 | 0 | 254 | 450 | 319 | 650 | 375 | 138 | 6 | 421 | |
| sim | 8.42 | 17.49 | 8.91 | 1.05 | 0.00 | 7.03 | 13.43 | 5.73 | 10.50 | 5.67 | 4.05 | 0.08 | 10.85 | |
| FN | 1029 | 778 | 984 | 1497 | 1530 | 1382 | 1045 | 1129 | 989 | 1135 | 1398 | 1506 | 1136 | |
| sim | 26.88 | 20.31 | 27.94 | 53.23 | 41.31 | 50.21 | 36.83 | 28.40 | 26.59 | 29.30 | 40.63 | 41.49 | 31.88 | |
| TN | 6629 | 6701 | 6597 | 6567 | 6560 | 6664 | 6617 | 6611 | 6793 | 6628 | 6583 | 6546 | 6680 | |
| Hit | 0.213 | 0.298 | 0.254 | 0.014 | 0.000 | 0.024 | 0.185 | 0.141 | 0.222 | 0.142 | 0.053 | 0.021 | 0.149 | |
| sim | 0.222 | 0.298 | 0.193 | 0.007 | 0.000 | 0.015 | 0.117 | 0.160 | 0.212 | 0.165 | 0.035 | 0.011 | 0.126 | |
| Error | 0.224 | 0.313 | 0.234 | 0.022 | 0.000 | 0.141 | 0.244 | 0.220 | 0.257 | 0.226 | 0.083 | 0.008 | 0.207 | |
| sim | 0.148 | 0.241 | 0.133 | 0.016 | 0.000 | 0.086 | 0.181 | 0.172 | 0.166 | 0.133 | 0.054 | 0.004 | 0.141 | |
| Under | 0.564 | 0.389 | 0.513 | 0.964 | 1.000 | 0.835 | 0.571 | 0.640 | 0.522 | 0.632 | 0.865 | 0.972 | 0.644 | |
| sim | 0.630 | 0.461 | 0.674 | 0.977 | 1.000 | 0.900 | 0.702 | 0.668 | 0.622 | 0.701 | 0.911 | 0.985 | 0.733 | |
| Over | 0.335 | 0.353 | 0.348 | 0.371 | 0.000 | 0.482 | 0.364 | 0.334 | 0.417 | 0.362 | 0.387 | 0.12 | 0.401 | |
| sim | 0.348 | 0.424 | 0.397 | 0.454 | 0.000 | 0.557 | 0.462 | 0.289 | 0.393 | 0.313 | 0.506 | 0.11 | 0.483 | |
| Score | 0.472 | 0.486 | 0.490 | 0.432 | 0.450 | 0.389 | 0.448 | 0.434 | 0.461 | 0.431 | 0.431 | 0.453 | 0.439 | |
| sim | 0.503 | 0.508 | 0.490 | 0.426 | 0.450 | 0.400 | 0.428 | 0.463 | 0.489 | 0.479 | 0.425 | 0.449 | 0.446 | |
[…]
| FP-sim | 15.83 | 29.93 | 15.62 | 1.76 | 0.00 | 12.58 | 24.30 | 14.17 | 18.94 | 12.00 | 6.84 | 0.27 | 17.76 | |
| FPne-sim | 7.20 | 12.38 | 6.58 | 0.70 | 0.00 | 5.27 | 10.94 | 8.38 | 8.37 | 6.25 | 2.70 | 0.19 | 6.81 | |
| FPun-sim | 8.63 | 17.54 | 9.03 | 1.07 | 0.00 | 7.31 | 13.36 | 5.80 | 10.57 | 5.75 | 4.14 | 0.08 | 10.95 | |
| FN-sim | 31.54 | 22.55 | 32.06 | 47.73 | 48.90 | 43.66 | 33.43 | 33.87 | 30.37 | 33.61 | 45.12 | 48.29 | 36.24 | |
| Hit | 0.155 | 0.239 | 0.189 | 0.010 | 0.000 | 0.016 | 0.137 | 0.100 | 0.165 | 0.106 | 0.036 | 0.015 | 0.105 | |
| sim | 0.154 | 0.240 | 0.174 | 0.009 | 0.000 | 0.014 | 0.125 | 0.100 | 0.155 | 0.103 | 0.033 | 0.012 | 0.102 | |
| Error | 0.159 | 0.261 | 0.178 | 0.015 | 0.000 | 0.110 | 0.192 | 0.165 | 0.192 | 0.162 | 0.059 | 0.005 | 0.147 | |
| sim | 0.134 | 0.229 | 0.147 | 0.013 | 0.000 | 0.094 | 0.170 | 0.144 | 0.164 | 0.129 | 0.051 | 0.004 | 0.127 | |
| Under | 0.403 | 0.268 | 0.373 | 0.627 | 0.647 | 0.563 | 0.390 | 0.447 | 0.375 | 0.450 | 0.574 | 0.635 | 0.449 | |
| sim | 0.429 | 0.299 | 0.415 | 0.629 | 0.647 | 0.580 | 0.425 | 0.467 | 0.407 | 0.475 | 0.586 | 0.639 | 0.471 | |
| Over | 0.183 | 0.315 | 0.227 | 0.023 | 0.000 | 0.171 | 0.224 | 0.163 | 0.266 | 0.165 | 0.086 | 0.004 | 0.206 | |
| sim | 0.183 | 0.320 | 0.230 | 0.023 | 0.000 | 0.169 | 0.229 | 0.159 | 0.264 | 0.150 | 0.089 | 0.005 | 0.211 | |
| Score | 0.485 | 0.486 | 0.493 | 0.466 | 0.468 | 0.428 | 0.461 | 0.453 | 0.474 | 0.458 | 0.461 | 0.474 | 0.461 | |
| sim | 0.493 | 0.498 | 0.496 | 0.466 | 0.468 | 0.432 | 0.462 | 0.461 | 0.478 | 0.469 | 0.462 | 0.473 | 0.466 |

### v2

@@
-CLEME2.0: Towards More Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction
+CLEME2.0: Towards Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction
 Abstract
-The paper focuses on improving the interpretability of Grammatical Error Correction (GEC) metrics, which receives little attention in previous studies.
-To bridge the gap, we propose CLEME2.0, a reference-based evaluation strategy that can describe four elementary dimensions of GEC systems, namely hit-correction, error-correction, under-correction, and over-correction.
-They collectively contribute to revealing the critical characteristics and locating drawbacks of GEC systems.
-Evaluating systems by Combining these dimensions leads to high human consistency over other reference-based and reference-less metrics.
-Extensive experiments on 2 human judgement datasets and 6 reference datasets demonstrate the effectiveness and robustness of our method.111All the codes will be released after the peer review.
-CLEME2.0: Towards More Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction
-Jingheng Ye1††thanks: indicates equal contribution., Zishan Xu1∗, Yinghui Li1, Xuxin Cheng2, Linlin Song3, Qingyu Zhou, Hai-Tao Zheng1,4††thanks: Corresponding author: Hai-Tao Zheng.
-(E-mail: zheng.haitao@sz.tsinghua.edu.cn), Ying Shen5, Xin Su6 1Tsinghua University, 2Peking University 3Huazhong University of Science and Technology 4Peng Cheng Laboratory, 5Sun-Yat Sen University, 6Tencent {yejh22,xzs23}@mails.tsinghua.edu.cn
+The paper focuses on the interpretability of Grammatical Error Correction (GEC) evaluation metrics, which received little attention in previous studies.
+To bridge the gap, we introduce CLEME2.0, a reference-based metric describing four fundamental aspects of GEC systems: hit-correction, wrong-correction, under-correction, and over-correction.
+They collectively contribute to exposing critical qualities and locating drawbacks of GEC systems.
+Evaluating systems by combining these aspects also leads to superior human consistency over other reference-based and reference-less metrics.
+Extensive experiments on two human judgment datasets and six reference datasets demonstrate the effectiveness and robustness of our method, achieving a new state-of-the-art result.
+Our codes are released at https://github.com/THUKElab/CLEME.
+CLEME2.0: Towards Interpretable Evaluation by Disentangling Edits
+for Grammatical Error Correction
+Jingheng Ye1††thanks: Equal Contribution., Zishan Xu111footnotemark: 1, Yinghui Li1, Linlin Song2, Qingyu Zhou3, Hai-Tao Zheng1,4††thanks: Corresponding Author: Hai-Tao Zheng.
+(E-mail: zheng.haitao@sz.tsinghua.edu.cn), Ying Shen5, Wenhao Jiang6, Hong-Gee Kim7, Ruitong Liu1, Xin Su8, Zifei Shan8 1Tsinghua University, 2Huazhong University of Science and Technology, 3ByteDance Inc., 4Peng Cheng Laboratory, 5Sun-Yat Sen University, 6Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ), 7Seoul National University, 8Tencent {yejh22,xzs23}@mails.tsinghua.edu.cn
 1 Introduction
-Grammatical Error Correction (GEC) is the task of automatically detecting and correcting all grammatical errors in a given text Li et al.
-(2022a, b); Ma et al.
-(2022); Ye et al.
+The task of Grammatical Error Correction (GEC) automatically detects and corrects grammatical errors in a given text Bryant et al.
 (2023).
-A core component of any NLP tasks is the development of automatic metrics that can objectively measure model performance Bryant et al.
+A core component of GEC is the development of automatic metrics that can objectively measure model performance Kobayashi et al.
+(2024b); Ye et al.
+(2023c).
+However, proposing appropriate evaluation of GEC has long been a challenging task Madnani et al.
+(2011), due to the subjectivity Bryant and Ng (2015), complexity Mita et al.
+(2019) and subtlety Choshen and Abend (2018) of GEC.
+Recent research efforts have focused on developing GEC metrics that closely align with human judgements Koyama et al.
+(2024), whereas the interpretability of these metrics has received less emphasis.
+We define the interpretability of metrics as their capacity to disclose concerned characteristics of systems, which is crucial for identifying weaknesses in a given GEC system.
+It is generally recognized that excellent systems should adhere to the principles of grammaticality and faithfulness Bryant et al.
 (2023).
-However, proposing appropriate evaluation of GEC has long been a challenging task Madnani et al.
+Grammaticality demands that all grammatical errors be accurately corrected, while faithfulness ensures that corrections retain the original meaning and syntactic structure.
+Nevertheless, the commonly utilized GEC metrics Bryant et al.
+(2017); Dahlmeier and Ng (2012a) are PRF-based (Precision, Recall, and F scores).
+We claim that PRF-based metrics fail to effectively capture subtle dimensions of GEC systems, consequently hindering progress.
+As illustrated in Figure 1, the edits [werewas] and [forin] in Hyp.
+1 are regarded as 2 FP + FN edits by ERRANT Bryant et al.
+(2017) or 2 FP edits by CLEME Ye et al.
+(2023c).
+Meanwhile, the edit [of] in Hyp.
+2 is categorized as an FP edit for both ERRANT and CLEME.
+Despite this, these two categories of FP edits carry different implications.
+The former type is correctly placed but wrongly modified, whereas the latter is incorrectly positioned.
+The inability to differentiate between these FP edits results in ambiguous interpretations of P/R/F0.5 scores, as they fail to quantify grammaticality and faithfulness.
+Thus, we introduce CLEME2.0, an interpretable reference-based metric describing four fundamental aspects of GEC systems: 1) the hit-correction score reflects the degree to which a system accurately corrects grammatical errors, 2) the wrong-correction score denotes the degree of incorrect corrections made, 3) the under-correction score reveals the degree of missing corrections, and 4) the over-correction score measures the degree of excessive corrections.
+An excellent GEC system should gain a higher hit-correction score and lower wrong-correction, under-correction, and over-correction scores.
+The initial three aspects assess grammaticality, whereas the over-correction score pertains to faithfulness, given that it often alters the original meaning, a challenge notably observed with LLMs Coyne et al.
 (2023); Li et al.
-(2023a), due to the subjectivity Bryant and Ng (2015), complexity Mita et al.
-(2019) and subtlety Choshen and Abend (2018) of GEC Napoles et al.
-(2015).
-Recent studies have been trying to develop GEC metrics that can achieve high correlations with human judgements Yoshimura et al.
-(2020a), with less attention paid to the interpretability of the automatic metrics.
-We define the interpretability of metrics as their ability to reveal the concerned characteristics of systems, which is vital in locating the drawbacks of a certain system.
-It is well-acknowledged that excellent GEC systems, which usually conform to the principle of minimal editing, should adhere to two gold principles, namely grammaticality and faithfulness.
-Grammaticality necessitates that all grammatical errors should be accurately corrected, while faithfulness requires that the corrections maintain the original textual meaning and syntactic structure.
-However, the widely-adopted mainstream GEC metrics Bryant et al.
-(2017); Ye et al.
-(2023c) indicate the GEC performance by precision, recall, and F scores, which can hardly characterize these critical dimensions of GEC systems, thus hindering the development.
-Therefore, we propose CLEME2.0, a more interpretable reference-based evaluation strategy that can describe four fundamental aspects of GEC systems: hit-correction, error-correction, under-correction, and over-correction.
-The first three aspects are responsible for describing grammaticality, while the last one is for faithfulness since the over-correction edits tend to change the original semantics, especially for LLMs Coyne et al.
-(2023).
-To achieve this, CLEME2.0 distinguishes between necessary and unnecessary corrections and disentangles edits into four main types: true positive (TP), necessary false positive (FP), unnecessary false positive (FP), false negative (FN) edits.222True negative edits are not considered in our method.
-For example in Figure 1, the Hyp.1 makes three necessary edits on the right positions, where [the ] is a TP edit but two of others ([were was] and [for in]) are FP edits since they are not covered in the reference.
-So Hyp.1 tends to mistakenly correct grammatically errors.
-On the other hand, Hyp.2 makes extra two FP edits ([ of] and ([century centuries]) since the reference does not correct the right positions, indicating the occurrence of two under-correction phenomena.
-Additionally, [for for] of the Hyp.2 is considered as an FN edit, which means the occurrence of an under-correction phenomenon.
-Since the edit disentanglement is based on the chunk partition technique proposed by CLEME, so we dub this strategy as CLEME2.0.
-Disentangling edits enables us to investigate concrete dimensions of GEC systems by computing upon an evaluation dataset four disentangled scores: hit-correction, error-correction, under-correction, and over-correction scores.
-In contrast to mainstream GEC metrics like ERRANT Bryant et al.
-(2017) and MaxMatch Dahlmeier and Ng (2012a) that reveal the system performance by P/R/F0.5, this disentanglement can provide an interpretable insight into fine-grained dimensions responsible for describing critical characteristics of GEC systems.
-Then, we integrate these disentangled scores into a comprehensive score using linear weighted summation, placing different emphases on disentangled scores.
-We leverage the comprehensive score to indicate the system performance from a global perspective.
-Similar to CLEME Ye et al.
-(2023c), CLEME2.0 also supports the evaluation based on either correction dependence or correction independence assumptions, providing a flexible option.
-Besides, we assume that edits with various extents of modification should affect distinctively the evaluation results.
-Therefore, we incorporate two edit weighting techniques into CLEME2.0, namely similarity-based weighting Gong et al.
+(2023a).
+To achieve this, CLEME2.0 first distinguishes between necessary and unnecessary false positive (FP) edits.
+The idea is that necessary FP edits indicate the system’s wrong-correction degree, while unnecessary FP edits reveal the system’s over-correction degree.
+As shown in the bottom block of Figure 1, [werewas] and [forin] in Hyp.
+1 are regarded as FP edits, while [of] in Hyp.
+2 is considered as an FP edit.
+As a result, CLEME2.0 establishes a one-to-one relationship between four distinct system aspects and four types of edits: hit-correction v.s.
+TP, wrong-correction v.s.
+FP, under-correction v.s.
+FN, and over-correction v.s.
+FP.
+Unlike conventional GEC metrics like ERRANT Bryant et al.
+(2017) and MaxMatch Dahlmeier and Ng (2012a) that evaluate using P/R/F0.5 scores, it offers a nuanced view into the detailed aspects necessary for characterizing critical features of GEC systems.
+These separated scores are then consolidated into an overall score via linear weighted summation, giving varying importance to these distinct scores.
+This aggregate score provides a holistic measure of system performance.
+Similar to CLEME, our method adopts the chunk partition technique and supports evaluations based on either correction dependence or correction independence assumptions, so we dub the metric as CLEME2.0.
+Moreover, we propose that edits of varying modification levels should uniquely influence the evaluation outcomes.
+For example, corrections involving punctuation are often less significant than corrections of content words.
+Therefore, we integrate two edit weighting techniques into CLEME2.0, similarity-based weighting Gong et al.
 (2022) and LLM-based weighting.
-Specifically, the techniques compute an important weight for each edit using a language model rather than treating each edit equally, thus equipping CLEME2.0 with abilities to capture context semantics and overcome the defect of traditional measures relying on superficial form similarity Kobayashi et al.
+In particular, these methods compute a specific weight for each edit through a language model rather than assigning equal weight to all edits, thereby enabling CLEME2.0 to grasp contextual semantics and address the limitations of conventional metrics that depend on surface-level form similarity Kobayashi et al.
 (2024a).
-To verify the effectiveness of CLEME2.0, we conduct extensive experiments on 2 human judgment datasets (GJG15 Grundkiewicz et al.
+To verify the effectiveness of CLEME2.0, we conduct extensive experiments on two human judgment datasets (GJG15 Grundkiewicz et al.
 (2015) and SEEDA Kobayashi et al.
 (2024b)), where our method consistently achieves high correlations.
-We also demonstrate the robustness of CLEME2.0 by computing the evaluation results based on 6 reference datasets with disparate annotation styles.
+We also demonstrate the robustness of CLEME2.0 by computing the evaluation results based on six reference datasets with disparate annotation styles.
 In summary, our contributions are three folds:
-We propose CLEME2.0, a more interpretable evaluation strategy, which is beneficial to reveal crucial characteristics of GEC systems.
-We boost CLEME2.0 with two edit weighting techniques, including similarity-based and LLM-based weighting, to overcome the inability of traditional reference-based metrics.
+We introduce CLEME2.0, an interpretable reference-based metric, which is beneficial to reveal crucial aspects of GEC systems.
+We enhance CLEME2.0 by incorporating two edit weighting techniques, addressing the limitations of conventional reference-based metrics in capturing semantics.
 Extensive experiments and analyses are conducted to confirm the effectiveness and robustness of our proposed method.
@@
 Reference-based metrics.
-Reference-based metrics evaluate GEC systems by referencing manually written materials Ye et al.
-(2023a, b); Huang et al.
+Reference-based metrics evaluate GEC systems by comparing their outputs to manually written references Ye et al.
+(2022, 2023a, 2023b); Huang et al.
 (2023); Li et al.
-(2024a).
-The M2 scorer Dahlmeier and Ng (2012b) identifies optimal edit sequences between source sentences and system hypotheses, using the F0.5 score.
+(2024c, 2022c, 2022b, d); Ma et al.
+(2022); Zhang et al.
+(2023, 2025a); Li et al.
+(2025c).
+The M2 scorer Dahlmeier and Ng (2012b) identifies optimal edit sequences between source sentences and system hypotheses, using F0.5 scores.
 However, this method can inflate scores by manipulating edit boundaries.
-Bryant et al.
-(2017) proposed ERRANT, which improves edit extraction with a linguistically-informed alignment algorithm, but it remains language-dependent and biased in multi-reference evaluation.
-Napoles et al.
-(2015) introduced GLEU, an n-gram-based metric inspired by BLEU for GEC evaluation.
-Ye et al.
-(2023c) proposed CLEME to eliminate bias in multi-reference evaluation by transforming the source, hypothesis, and references into chunk sequences with consistent boundaries, providing unbiased F0.5 scores.
-Gong et al.
-(2022) introduce PT-M2, focusing on scoring changed words extracted by the M2 metric.
+To mitigate this problem, ERRANT Bryant et al.
+(2017) improves edit extraction through a linguistically-informed alignment algorithm, but it remains language-dependent and biased in multi-reference evaluation.
+CLEME Ye et al.
+(2023c) further provides unbiased F0.5 scores and introduces an extra correction assumption for multi-reference evaluation.
+PT-M2 Gong et al.
+(2022) combines PT-based and existing GEC metrics for higher correlations with human judgments.
 Reference-less metrics.
-To overcome the limitations of reference-based metrics, recent research focus on reference-less scoring.
+To overcome the limitations of reference-based metrics, recent studies focus on reference-less scoring.
 Inspired by quality estimation in NMT Liu et al.
@@
 Yoshimura et al.
-(2020b) introduce SOME, which uses sub-metrics optimized for manual assessment with regression models.
+(2020) introduce SOME, which uses sub-metrics optimized for manual assessment with regression models.
 Scribendi Score Islam and Magnani (2021) combines language perplexity and token/Levenshtein distance ratios.
@@
 3.1 Edit Extraction
-The first step is edit extraction.
-Given a source sentence and a target (either hypothesis or reference) sentence , this step is to extract the edits describing the modification from to .
+Given a source sentence and a target (either hypothesis or reference) sentence , we extract the edits describing the modification from to .
 Here, we utilize the chunk partition technique from CLEME Ye et al.
@@
 3.2 Disentangled Scores
-For the purpose of computing disentangled scores, we initially disentangle edits into four core types.
+To compute disentangled scores, we initially disentangle edits into four elementary types.
 1) TP edits refer to the corrected/dummy hypothesis chunks that share the same tokens as the corresponding reference chunks.
@@
 (2017); Li et al.
-(2023b) do not distinguish between FP and FP, treating both as FP, thereby resulting in confusion between error-correction and over-correction.
+(2023c) do not distinguish between FP and FP, treating both as FP, thereby resulting in confusion between wrong-correction and over-correction.
 Actually, we have .
@@
 The formula is as follows:
-Error-correction score.
-Conversely, the error-correction score is defined as the ratio of FP edits to all necessary reference edits.
+Wrong-correction score.
+Conversely, the wrong-correction score is defined as the ratio of FP edits to all necessary reference edits.
 This score seeks to evaluate the degree to which systems generate erroneous corrections for grammatical errors.
@@
 Under-correction score.
-Similarly, the under-correction score is proposed to measure the degree to which systems omit to correct grammatical errors, which is computed as follow:
+Similarly, the under-correction score is proposed to measure the degree to which systems omit to correct grammatical errors, which is computed as follows:
 Over-correction score.
@@
 This score is determined by the proportion of FP edits to all hypothesis corrected/dummy edits, aiming to gauge the level to which systems offer excessive corrections:
+With the disentangled scores indicating disparate aspects of GEC systems, researchers can identify specific weaknesses and implement targeted improvements without expensive human labor.
 3.3 Comprehensive Score
@@
 These weights are then incorporated into the calculation of the aforementioned disentangled scores as depicted in Equation (1) (4).
-Take the hit-correction score as a typical example, we reformulate the Equation (1) as follow:
+Take the hit-correction score as a typical example, we reformulate the Equation (1) as follows:
 Similarity-based weighting.
-We use PTScore from Gong et al.
-(2022) to provide edit weights.
-Through simulating a partially accurate version of the source sentence , PTScore can assign individual weights to edits, in spite of multiple edits in a sentence.
+We use PTScore to assign edit weights Gong et al.
+(2022).
 Since it performs based on BERTScore Zhang et al.
-(2019) designed to compute similarity scores for text generation, we call this technique as similarity-based weighting.
-The computation process is as follows:
-where the function is intended to replace a specific chunk of the source with the corrected/dummy hypothesis chunk .
-Here, denotes the reference sentence.
-Comprehensive details can be found in Gong et al.
+(2019), a tool for evaluating text generation through similarity scores, we refer to this technique as similarity-based weighting.
+The rationale is to prioritize edits with a more significant modification of the meaning and quality of the text.
+By simulating a partially accurate version of the source sentence , PTScore can associate specific weights to edits within a sentence. The computation process is as follows:
+where is the reference sentence, while the function is used to replace a specific chunk of the source with the corrected/dummy hypothesis chunk .
+A positive weight indicates a beneficial correction, whereas a negative value suggests a wrong correction.
+The absolute value is utilized as the edit weight following (Gong et al., 2022), and the significance111For more detailed analysis, refer to our case study in Section 5.1 and PT-M2 Gong et al.
 (2022).
+of an edit in a sentence grows with a larger .
 LLM-based weighting.
-In light of the impressive semantic understanding capabilities of LLMs Tan et al.
-(2023); Li et al.
-(2024b); Yu et al.
-(2024); Huang et al.
-(2024); Du et al.
-(2024); Xu et al.
-(2024); Qin et al.
-(2024b), some recent work has sought their use in evaluating assorted NLP tasks Pavlovic and Poesio (2024); Chen et al.
-(2024); Qin et al.
-(2024a), GEC evaluations included Sottana et al.
+Recent studies have begun investigating the effectiveness of LLM-based evaluation, known for their advanced semantic comprehension Qin et al.
+(2024b), in assessing various NLP tasks Pavlovic and Poesio (2024); Sottana et al.
 (2023).
-Drawing inspiration from this trend, we employ Llama-2-7B Touvron et al.
-(2023) as a scorer to acquire the importance score for each edit.
-These scores range from 1 to 5, with higher scores indicating increased edit importance.
-We provide the implementation details and the prompting template in Appendix D.
+Building on this trend, we prompt Llama-2-7B Touvron et al.
+(2023) to assign edit weights from 1 to 5, where a higher value signifies more critical edits.
+This methodology is rooted in the idea that LLMs, due to their extensive training on diverse data, are adept at grasping intricate language patterns and text structure.
+Detailed implementation instructions and the prompting framework are available in Appendix A.
 4 Experiments
@@
 (2016).
-Notably, SN-8GEC is partitioned into 4 sub-sets, namely E-Minimal, E-Fluency, NE-Minimal, and NE-Fluency.
-A more thorough breakdown of these datasets and the statistics are provided in Appendix A.
+Notably, SN-8GEC is partitioned into 4 sub-sets, i.e., Expert-Minimal, Expert-Fluency, Non-Expert-Minimal, and Non-Expert-Fluency.
+A more thorough breakdown of these datasets and the statistics is provided in Appendix B.
 Corpus and sentence levels.
@@
 Given the metric , source sentences , hypothesis sentences and reference sentences , 1) corpus-level metrics compute the system score based on the whole corpus , and 2) sentence-level metrics use the average of the sentence-level scores .
+Trade-off factors.
+We leverage a cross-evaluation search method to identify two optimal sets of trade-off factors for both the corpus and sentence levels.
+At the corpus level, the factors are assigned as , while for the sentence level, they are adjusted to .
+The details of the chosen values of trade-off factors can be seen in Appendix B.5.
+Evaluation assumptions.
+CLEME can evaluate GEC systems based on correction dependence (-dep) or independence (-ind) assumptions.
+The correction independence assumption offers a more relaxed edit-matching process, implying that systems might yield better scores when multiple references are available.
+Inspired by this work, CLEME2.0 also supports both assumptions, and we will study their effects on our method.
+4.2 Results of GJG15 Ranking
+The correlations between the GEC metrics and human judgments on the GJG15 rankings are shown in Table 1, and we have the following insights.
+CLEME2.0 outperforms other metrics at both corpus and sentence levels.
+For corpus-level, CLEME2.0-sim-ind achieves the highest average correlations, closely followed by CLEME2.0-sim-dep.
+CLEME2.0-ind and CLEME2.0-dep can also gain comparable correlations with other metrics, even though they do not utilize any edit weighting techniques.
+On the other hand, sentence-level metrics exhibit a similar pattern.
+SentCLEME2.0-sim-dep and SentCLEME2.0-sim-ind achieve the highest Pearson and Spearson correlations, respectively.
+These results significantly demonstrate the effectiveness and robustness of our proposed method across different settings.
+Sentence-level metrics outperform their corpus-level counterparts.
+This observation is consistent with recent studies Gong et al.
+(2022); Ye et al.
+(2023c).
+This is because system-level rankings treat each sample equally regardless of edit numbers, mirroring how sentence-level metrics are evaluated.
+On the other hand, corpus-level metrics emphasize samples with more edits, thus causing the gap between automatic metrics and human evaluation.
+SentPT-M2 shows superior performance on CoNLL-2014 but performs worse on BN-10GEC, E-Minimal, and NE-Fluency compared to our approach, revealing a lack of robustness of the metric.
+Generally, our method aligns more closely with human assessments than existing popular metrics.
+Notably, our method with similarity-based weighting surpasses unweighted ones, thanks to the integration of semantic factors.
+However, on E-Minimal and NE-Minimal, weighted and unweighted results are comparable.
+We suspect this is because these datasets have minimal yet crucial annotations, reducing the possibility of varying weights and the efficacy of edit weighting.
+Furthermore, we present comprehensive results of CLEME2.0 on CoNLL-2014 and offer insights into our method for analyzing and identifying weaknesses in GEC systems in Appendix C.5.
 | Metric | SEEDA-S | SEEDA-E | Avg. | ||
@@
 | SentCLEME2.0-sim-ind (Ours) | 0.921 | 0.886 | 0.957 | 0.960 | 0.931 |
-| Metric | EW | TS | Avg. | ||
-| CLEME2.0-dep-Hit | 0.599 | 0.593 | 0.673 | 0.648 | 0.628 |
-| CLEME2.0-dep-Error | -0.444 | -0.533 | -0.526 | -0.593 | -0.524 |
-| CLEME2.0-dep-Under | 0.496 | 0.599 | 0.576 | 0.659 | 0.583 |
-| CLEME2.0-dep-Over | 0.118 | 0.269 | 0.073 | 0.275 | 0.253 |
-| SentCLEME2.0-dep-Hit | 0.594 | 0.593 | 0.672 | 0.648 | 0.627 |
-| SentCLEME2.0-dep-Error | -0.405 | -0.429 | -0.489 | -0.500 | -0.456 |
-| SentCLEME2.0-dep-Under | 0.489 | 0.511 | 0.572 | 0.582 | 0.539 |
-| SentCLEME2.0-dep-Over | -0.247 | -0.363 | -0.346 | -0.440 | -0.349 |
-Trade-off factors.
-We employ a cross-evaluation approach to determine two optimal configurations for trade-off factors applicable at the corpus and sentence levels, respectively.
-At the corpus level, we assign the factors as .
-Conversely, at the sentence level, these are adjusted to .
-Additionally, CLEME2.0 metrics named with “sim” are based on similarity-based edit weighting, and we leave the analysis of LLM-based edit weighting to Section 5.2.
-Evaluation Assumptions.
-Ye et al.
-(2023c) propose evaluating systems based on one of two assumptions, namely correction dependence and correction independence.
-In short, the correction independence assumption offers a more relaxed edit matching process, implying that systems are more inclined to yield higher scores when multiple references are accessible.
-Inspired by this work, CLEME2.0 also supports both assumptions, and we will study their impact on our method.
-4.2 Results of GJG15 Ranking
-The correlations between the GEC metrics and human judgments on the GJG15 rankings are shown in Table 1, and we have the following insights.
-CLEME2.0 outperforms other metrics at both corpus and sentence levels.
-For corpus-level, CLEME2.0-sim-ind achieves the highest average correlations, closely followed by CLEME2.0-sim-dep.
-CLEME2.0-ind and CLEME2.0-dep can also gain comparable correlations with other metrics, in spite of the fact that they do not utilize any edit weighting techniques.
-On the other hand, sentence-level metrics exhibit a similar pattern.
-SentCLEME2.0-sim-dep and SentCLEME2.0-sim-ind achieve the highest Pearson and Spearson correlations, respectively.
-These results significantly demonstrate the effectiveness and robustness of our proposed method across different settings.
-Sentence-level metrics outperform their corpus-level counterparts.
-This observation aligns with recent studies Gong et al.
-(2022); Ye et al.
-(2023c).
-This is because system-level rankings treat each sample equally, which is consistent with the evaluation process of sentence-level metrics.
-In contrast, corpus-level metrics allow samples with more edits to significantly influence the results.
-SentPT-M2 exhibits the best performance on CoNLL-2014, but its results on BN-10GEC, E-Minimal, and NE-Fluency are inferior compared to our method.
-In general, our method aligns more consistently with human judgments than existing mainstream metrics.
-Particularly, most weighted outcomes outshine the unweighted ones, attributable to the incorporation of semantic considerations.
-However, on E-Minimal and NE-Minimal, the unweighted and weighted results exhibit comparability.
-We conjecture that this could be due to the annotations in these datasets being minimal yet decisive, reducing the possibility of generating diverse weights.
 4.3 Results of SEEDA Ranking
-We conduct a supplementary experiment on the SEEDA-Sentence and SEEDA-Edit datasets, comparing our method with a wide range of GEC metrics.
-Table 2 demonstrates again that our approach obtains the optimal results on both datasets.
-Kobayashi et al.
-(2024b) claim that the correlations of most metrics tend to decline when shifting from classical to neural systems in evaluation.
-This suggests that traditional metrics may struggle when assessing more extensively edited and fluent corrections generated by neural systems.
-However, our method is still able to address these challenges and obtain even better results.
-The results on SEEDA-Edit surpass those on SEEDA-Sentence due to the finer granularity of SEED-Edit, which is more consistent with the functioning of CLEME2.0.
+We carry out an additional experiment on the SEEDA-Sentence and SEEDA-Edit datasets, where we compare our method against various GEC metrics.
+As presented in Table 2, our approach consistently achieves the best outcomes across both datasets.
+According to Kobayashi et al.
+(2024b), the correlations of most metrics tend to decrease when transitioning from classical to neural evaluation systems.
+This implies that conventional metrics might face difficulties in evaluating the more extensively edited or fluent corrections produced by state-of-the-art neural GEC systems.
+Nevertheless, our method effectively tackles these challenges, delivering even improved performance for all indicators.
+The results for SEEDA-Edit exceed those for SEEDA-Sentence, due to the greater detail in SEEDA-Edit, aligning more closely with the operation of CLEME2.0.
 It is crucial to mention that reference-less metrics such as SOME and IMPARA yield high outcomes, in part, because these are fine-tuned on GEC data.
@@
 However, only moderate correlations are observable on SEEDA-Edit.
+| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | |
+| Source | Do one | who | suffered | from this disease keep it a secret | of infrom | their relatives ? |
+| Reference | Does one | who | suffers | from this disease keep it a secret | or inform | their relatives ? |
+| Hypothesis | Do one (0.028) | who | suffer (0.011) | from this disease keep it a secret | to inform (0.094) | their relatives ? |
+| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | Chunk 7 | Chunk 8 | Chunk 9 | |
+| Source | When we are | diagonosed out | with certain genetic | disease | , should we disclose | this result | to | our | relatives ? |
+| Ref. | When we are | diagnosed | with certain genetic | diseases | , should we disclose | this result | to | our | relatives ? |
+| Hyp. | When we are | diagnosed out (0.056) | with certain genetic | diseases (0.006) | , should we disclose | the results (0.019) | to | their (0.021) | relatives ? |
+| Dataset | Corpus-EW | Corpus-TS | Sentence-EW | Sentence-TS | ||||
+| CoNLL-2014 | 0.697 | 0.659 | 0.759 | 0.720 | 0.626 | 0.654 | 0.696 | 0.698 |
+| BN-10GEC | 0.732 | 0.764 | 0.796 | 0.813 | 0.638 | 0.637 | 0.708 | 0.698 |
+| E-Minimal | 0.709 | 0.786 | 0.779 | 0.819 | 0.642 | 0.692 | 0.715 | 0.747 |
+| E-Fluency | 0.760 | 0.786 | 0.831 | 0.841 | 0.642 | 0.665 | 0.720 | 0.714 |
+| NE-Minimal | 0.777 | 0.823 | 0.839 | 0.861 | 0.654 | 0.747 | 0.723 | 0.791 |
+| NE-Fluency | 0.823 | 0.692 | 0.849 | 0.709 | 0.664 | 0.791 | 0.742 | 0.830 |
+4.4 Results of LLM-based Weighting
+Table 4 presents the outcomes of LLM-based weighting, noting that its effectiveness is less favorable than similarity-based weighting.
+A likely reason is the coarse grading method of LLMs, which allocates edit weights from 1 to 5, unlike the finer continuous scale .
+Although Kobayashi et al.
+(2024a) argue that LLMs serve as effective evaluators for GEC, their research pertains to huge closed-source LLMs (GPT-4 and GPT-3.5) and involves specific prompt engineering.
+They also identify the importance of the LLM scale since GPT-3.5 may even obtain negative correlations with human judgments.
+In contrast, we employ a more straightforward approach with open-source LLama-2-7B.
+5 Analysis
+5.1 Case Study
+Table 3 demonstrates instances of CLEME2.0.
+In the first set, Chunks 3 and 5 are FPne edits contributing to the wrong-correction score, with a higher edit weight of Chunk 5 than Chunk 3 since Chunk 5 introduces an error that entirely alters the sentence’s meaning.
+In the second set, Chunk 2 obtains the highest edit weight of 0.056, underscoring its substantial influence on the evaluation.
+Despite the correct modification of “diagnosed”, the misuse of “out” remains, keeping the correction wrong.
+Chunk 4 illustrates a singular-to-plural correction in the source sentence, with a low weight indicating a minor impact.
+Chunks 6 and 8 showcase over-corrections.
+Chunk 6 leaves the original meaning unchanged, whereas Chunk 8 introduces a significant error by misusing a personal pronoun.
+The cases highlight the effectiveness of the weighting technique.
+Otherwise, all edits are given equal weight, failing to distinguish hypothesis edits with varying correction levels.
+We display the cases of LLM-based weighting in Appendix C.2.
+5.2 Ablation Study
+| Metric | EW | TS | Avg. | ||
+| CLEME2.0-dep-Hit | 0.599 | 0.593 | 0.673 | 0.648 | 0.628 |
+| CLEME2.0-dep-Wrong | -0.444 | -0.533 | -0.526 | -0.593 | -0.524 |
+| CLEME2.0-dep-Under | 0.496 | 0.599 | 0.576 | 0.659 | 0.583 |
+| CLEME2.0-dep-Over | 0.118 | 0.269 | 0.073 | 0.275 | 0.253 |
+| SentCLEME2.0-dep-Hit | 0.594 | 0.593 | 0.672 | 0.648 | 0.627 |
+| SentCLEME2.0-dep-Wrong | -0.405 | -0.429 | -0.489 | -0.500 | -0.456 |
+| SentCLEME2.0-dep-Under | 0.489 | 0.511 | 0.572 | 0.582 | 0.539 |
+| SentCLEME2.0-dep-Over | -0.247 | -0.363 | -0.346 | -0.440 | -0.349 |
+We conduct ablation studies on (Sent)CLEME2.0-dep to analyze the performance of individual disentangled scores.
+A preferable system has reduced wrong-correction, under-correction, and over-correction scores, so we report corrections between 1- with human judgments where is one of the scores.
+The outcomes are detailed in Table 5.
+Hit-correction and under-correction show moderate correlations.
+Over-correction scores have small positive correlations at the corpus level, with minimal negative correlations at the sentence level.
+Notably, wrong-correction scores display negative correlations, but this does not mean they do not impact the overall score.
+In reality, the trade-off factor for wrong-correction scores is relatively substantial.
+The hypothesis is that focusing evaluations only on wrong-correction scores might prefer systems that make only highly confident edits, potentially leading to assessment bias.
+Additionally, we utilize the similarity-weighting approach on CLEME to evaluate its efficacy, with the outcomes detailed in Appendix C.3.
+To examine our method on a broad scale, we also provide the average correlations obtained from a comprehensive analysis of all potential parameter settings.
+The results are found in Appendix C.4.
+| Metric | Time (Seconds) |
+| ERRANT | 33.4 |
+| GLEU | 21.5 |
+| CLEME-dep | 54.1 |
+| CLEME-ind | 54.1 |
+| (Sent)CLEME2.0-dep | 54.1 |
+| (Sent)CLEME2.0-ind | 54.1 |
+| (Sent)CLEME2.0-sim-ind | 88.4 |
+| (Sent)CLEME2.0-sim-ind | 87.6 |
+5.3 Efficiency
+This section provides a comparative analysis of the efficiency of our methods against other prevailing metrics.
+The experiments were executed on a GPU 3090 within the CoNLL-2014 framework, with the evaluation times of the AMU system reported.
+Our observations are as follows: (1) For ERRANT, the primary time expenditure is associated with edit extraction, lasting 33.4 seconds.
+(2) CLEME and CLEME2.0 primarily incur time costs from edit extraction at 33.4 seconds and chunk partitioning at 20.7 seconds.
+(3) For CLEME2.0-sim, the most significant time costs are assignable to edit extraction (33.4 s), chunk partitioning (20.7 s), and edit weighting (34.3 s).
+PT-M2 exhibits the slowest runtime when replicating existing mainstream methods, with its evaluation process taking several hours; thus, we did not report a precise runtime due to the time constraints.
+Some technical solutions can mitigate the runtime when evaluating a system using these metrics concurrently.
+For instance, when assessing a system with ERRANT, CLEME, and CLEME2.0, the minimum cumulative duration is calculated as 33.4 seconds for edit extraction, 20.7 seconds for chunk partitioning, and 34.3 seconds for edit weighting, totaling 88.4 seconds.
+6 Conclusion
+This paper introduces CLEME2.0, an interpretable evaluation metric for GEC that effectively highlights four key aspects of systems.
+By incorporating edit weighting techniques, we overcome the challenges traditional reference-based metrics face in recognizing semantic subtleties.
+Extensive experiments and analyses confirm the effectiveness and robustness of our method.
+We anticipate that CLEME2.0 will offer a valuable perspective in the GEC community.
+Limitation
+Limitation in languages and datasets.
+While CLEME2.0 is adaptable to various languages, its efficiency beyond English remains unverified.
+Additionally, the reference sets employed in our experiments stem from the CoNLL-2014 shared task, which involves a second language dataset.
+To confirm the robustness of our methods, it’s necessary to conduct further experiments using evaluation datasets that cover a range of languages and text domains.
+Finally, we highly encourage the creation of new GEC evaluation datasets to foster progress.
+Lack of further human evaluation for interpretability.
+The experiments discussed in the paper are primarily concerned with assessing the correlation between automatic metrics and human judgments.
+However, they fall short of providing a thorough analysis of the method’s interpretability.
+Although we showcase the strong correlation performance of CLEME2.0, its interpretability is still unverified.
+In future work, we will conduct human evaluation experiments to showcase the interpretability of our method.
+Ethics Statement
+In this paper, we validate the effectiveness and robustness of our proposed approach using the CoNLL-2014, BN-10GEC, and SN-8GEC reference datasets.
+These datasets are sourced from publicly available resources on legitimate websites and do not contain any sensitive data.
+Additionally, all the baselines employed in our experiments are publicly accessible GEC metrics, and we have duly cited the respective authors.
+We confirm that all datasets and baselines utilized in our experiments are consistent with their intended purposes.
+Acknowledgements
+This research is supported by National Natural Science Foundation of China (Grant No.
+62276154), Research Center for Computer Network (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No.
+2023A1515012914 and 440300241033100801770), Basic Research Fund of Shenzhen City (Grant No.
+JCYJ20210324120012033, JCYJ20240813112009013 and GJHZ20240218113603006), the Major Key Project of PCL (NO.
+PCL2024A08).
+Appendix A LLM-based Edit Weighting
+Because of the powerful semantic comprehension abilities of LLMs Qin et al.
+(2024a); Tan et al.
+(2024); Ye et al.
+(2025a); Yu et al.
+(2024a); Tang et al.
+(2025); Yan et al.
+(2025); Li et al.
+(2024e, 2022a); Du et al.
+(2024); Huang et al.
+(2024); Li et al.
+(2023b); Yu et al.
+(2024b); Li et al.
+(2025a); Kuang et al.
+(2024), recent studies Chu et al.
+(2025); Ye et al.
+(2025c, b, 2024); Hu et al.
+(2024); Chen et al.
+(2024); Su et al.
+(2025); Zou et al.
+(2025); Xu et al.
+(2025); Li et al.
+(2025b, 2024b); Zhang et al.
+(2025b); Li et al.
+(2024a) have generated interest in employing LLMs for text assessment on various NLP tasks.
+Building on this idea, we use Llama-2-7B Touvron et al.
+(2023) as a scorer to determine edit weights.
+The prompt for edit weighting is presented in Figure 3.
+We set the temperature to 0.1 to ensure consistent and certain results.
+We instruct the LLM to evaluate each edit individually to prevent interference from other grammatical errors.
+Edit weights vary from 1 to 5, with higher values representing a greater need for correction.
+We do not specify the types of edits to the LLM; instead, we allow the LLM to directly evaluate the importance of edits through its inherent language understanding abilities.
+An input is composed of an uncorrected sentence and a certain edit.
+Appendix B Details about GEC Meta-Evaluation
+B.1 Human Rankings
+GJG15 ranking.
+Grundkiewicz et al.
+(2015) propose the first large-scale human judgement dataset for 12 participating systems of the CoNLL-2014 shared task.
+In this assessment, 8 native speakers are asked to rank the systems’ outputs from best to worst.
+Two system ranking lists are generated using Expected Wins (EW) and TrueSkill (TS), respectively.
+SEEDA ranking.
+Kobayashi et al.
+(2024b) identify several limitations of the GJG15 ranking dataset, and propose a new human ranking dataset called SEEDA.
+SEEDA consists of corrections with human ratings along two different granularities: edit-based and sentence-based, covering 12 state-of-the-art systems, including large language models (LLMs), and two human corrections with different focuses.
+Three native English speakers participate in the annotation process.
+Similar to Grundkiewicz et al.
+(2015), the overall human rankings are derived from TrueSkill (TS) and Expected Wins (EW) based on pairwise judgments.
+B.2 Ranking Algorithms
+Our employed human judgments are originally pairwise comparisons, i.e., humans choose the better of two available system outputs.
+The overall rankings are derived by using ranking algorithms, including Expected Wins (EW) and TrueSkill (TS).
+Expected Wins (EW)
+EW Bojar et al.
+(2013) is a derived ranking metric that quantifies the theoretical number of wins a participant is expected to achieve against a defined set of opponents.
+It is calculated by summing the probability of winning against each opponent, where these probabilities are typically derived from an existing skill rating system.
+EW provides a single aggregate score for ranking, useful for pre-match seeding or assessing theoretical group performance.
+TrueSkill (TS)
+TS Sakaguchi et al.
+(2014) is a Bayesian skill rating system developed by Microsoft Research.
+Unlike simpler systems, TS models a participant’s skill as a probability distribution (), where represents the estimated skill level and quantifies the uncertainty in that estimate.
+Upon match outcomes, TS updates these distributions using Bayesian inference, allowing for rapid adjustments and robust ranking.
+A key advantage is its inherent support for multi-player or team-based matches and the explicit handling of draws.
+Participants are typically ranked by a conservative estimate of their skill, such as , which accounts for confidence.
+| Item | CoNLL-2014 | BN-10GEC | E-Minimal | E-Fluency | NE-Minimal | NE-Fluency |
+| # Sentence (Length) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) |
+| # Reference (Length) | 2,624 (22.8) | 13,120 (22.9) | 2,624 (23.2) | 2,624 (22.8) | 2,624 (23.0) | 2,624 (22.2) |
+| # Edit (Length) | 5,937 (1.0) | 36,677 (1.0) | 4,500 (1.0) | 8,373 (1.1) | 4,964 (0.9) | 11,033 (1.2) |
+| # Unchanged Chunk (Length) | 11,174 (4.8) | 93,496 (2.5) | 8,887 (6.3) | 12,823 (3.8) | 10,748 (5.1) | 14,086 (2.9) |
+| # Corrected/Dummy Chunk (Length) | 4,994 (1.3) | 26,948 (2.4) | 3,963 (1.2) | 6,305 (1.7) | 4,221 (1.2) | 6,892 (2.6) |
+| Metric | CoNLL-2014 | BN-10GEC | E-Minimal | E-Fluency | NE-Minimal | NE-Fluency | Avg. | |||||||
+| EW | TS | EW | TS | EW | TS | EW | TS | EW | TS | EW | TS | |||
+| SentGLEU | 0.784 | 0.828 | 0.756 | 0.826 | 0.624 | 0.581 | 0.785 | 0.846 | 0.218 | 0.142 | 0.778 | 0.848 | 0.668 ( 0.120) | |
+| 0.720 | 0.775 | 0.769 | 0.824 | 0.599 | 0.593 | 0.791 | 0.846 | 0.220 | 0.170 | 0.768 | 0.846 | 0.660 ( 0.131) | ||
+[…]
+| 0.742 | 0.747 | 0.786 | 0.830 | 0.626 | 0.588 | 0.813 | 0.764 | -0.003 | -0.137 | 0.830 | 0.857 | 0.620 ( 0.179) | ||
+[…]
+| 0.824 | 0.808 | 0.835 | 0.874 | 0.451 | 0.412 | 0.824 | 0.863 | -0.181 | -0.247 | 0.791 | 0.846 | 0.592 ( 0.233) | ||
+[…]
+| 0.725 | 0.758 | 0.209 | 0.143 | 0.368 | 0.335 | 0.808 | 0.846 | -0.167 | -0.247 | 0.762 | 0.825 | 0.447 ( 0.335) | ||
+[…]
+| 0.714 | 0.725 | 0.681 | 0.747 | 0.846 | 0.852 | 0.824 | 0.901 | 0.368 | 0.352 | 0.720 | 0.791 | 0.710 ( 0.091) | ||
+[…]
+| 0.709 | 0.720 | 0.681 | 0.747 | 0.819 | 0.835 | 0.857 | 0.923 | 0.423 | 0.401 | 0.654 | 0.720 | 0.707 ( 0.086) | ||
+[…]
+| 0.907 | 0.912 | 0.808 | 0.863 | 0.808 | 0.813 | 0.885 | 0.945 | 0.527 | 0.505 | 0.896 | 0.940 | 0.817 ( 0.051) | ||
+[…]
+| 0.868 | 0.879 | 0.753 | 0.824 | 0.808 | 0.841 | 0.879 | 0.956 | 0.544 | 0.527 | 0.835 | 0.923 | 0.803 ( 0.052) | ||
+B.3 Statistics of Reference Datasets
+Table 7 presents the statistics of all the reference sets involved in our experiments.
+B.4 Baseline Metrics
+In our evaluation, we compare our method with the following reference-based baseline metrics, including corpus and sentence-level variants:
+M2 and SentM2 Dahlmeier and Ng (2012a) dynamically extract the hypothesis edits with the maximum overlap of gold annotations by utilizing the Levenshtein algorithm.
+GLEU and SentGLEU Napoles et al.
+(2015) are BLEU-like GEC metrics based on n-gram matching, rewarding hypothesis n-grams that align with the reference but not the source, while penalizing those aligning solely with the source.
+GLEU is the main metric in JFLEG, an English GEC dataset that highlights holistic fluency edits.
+ERRANT and SentERRANT Bryant et al.
+(2017) are among the most widely recognized in grammatical error correction.
+They enhance the accuracy of edit extraction by employing a linguistically refined version of the Damerau-Levenshtein algorithm.
+PT-M2 and SentPT-M2 Gong et al.
+(2022) leverage pre-trained language model (PLM) to evaluate GEC systems.
+The main idea is similar to M2 and ERRANT, but they can leverage the knowledge of pre-trained language models to score edits effectively.
+CLEME and SentCLEME Ye et al.
+(2023c) are proposed to provide unbiased scores for multi-reference evaluation.
+Furthermore, the authors present the correction independence assumption, enabling CLEME to function under either the traditional correction dependence or correction independence assumptions.
+| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | |
+| Source | Do one | who | suffered | from this disease keep it a secret | of infrom | their relatives ? |
+| Reference | Does one | who | suffers | from this disease keep it a secret | or inform | their relatives ? |
+| Hypothesis | Do one (5) | who | suffer (5) | from this disease keep it a secret | to inform (1) | their relatives ? |
+| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | Chunk 7 | Chunk 8 | Chunk 9 | |
+| Source | When we are | diagonosed out | with certain genetic | disease | , should we disclose | this result | to | our | relatives ? |
+| Ref. | When we are | diagnosed | with certain genetic | diseases | , should we disclose | this result | to | our | relatives ? |
+| Hyp. | When we are | diagnosed out (5) | with certain genetic | diseases (5) | , should we disclose | the results (4) | to | their (5) | relatives ? |
+For the evaluation on SEEDA, we add extra evaluation metrics following the evaluation methods reported in Kobayashi et al. (2024b):
+GoToScorer Gotou et al.
+(2020): takes into account the difficulty of error correction when calculating the evaluation score.
+The difficulty is calculated based on the number of systems that can correct errors.
+SOME Yoshimura et al. (2020): optimizes human evaluation by fine-tuning BERT separately for criteria such as grammaticality, fluency, and meaning preservation.
+IMPARA Maeda et al. (2022): incorporates a quality assessment model fine-tuned using BERT parallel data and a similarity model that takes into account the effects of editing.
+B.5 Details of Determining Trade-off Factors
+A cross-validation approach was employed on the six reference sets of GJG15 to determine the optimal set.
+Five of the six reference sets were selected, and an exhaustive exploration of all trade-off factors was conducted.
+The candidate factors were evaluated at intervals determined by a grid value of 0.05.
+The optimal factors were then identified and applied to the remaining reference set, yielding resultant corrections.
+We reiterated this process six times to ascertain the final set of trade-off factors, which exhibited the highest average correction for the remaining reference sets.
+Appendix C Extra Results
+C.1 Results of Full References
+The results without excluding unchanged reference sentences are presented in Table 8.
+We observe an obvious performance reduction in traditional metrics, especially in NE-Minimal, which contains numerous under-corrections due to annotation by non-experts under the minimal editing guideline.
+We remove 470 unchanged references in E-Minimal and 612 unchanged references in NE-Minimal.
+In particular, SentERRANT, SentCLEME-dep, and SentCLEME-ind exhibit negative correlations in NE-Minimal, revealing their lack of robustness.
+Many metrics also undergo a significant decrease in E-Minimal except CLEME2.0.
+In the case of E-Minimal, many metrics also show a marked decline, except for CLEME2.0.
+Our approach achieves the highest or comparable correlations across all reference sets, underscoring its robustness.
+C.2 Case Study of LLM-based Weighting
+In Table 9, we report instances of CLEME2.0 using LLM-based weighting.
+We notice distinct preferences when comparing similarity-based and LLM-based weighting methods.
+In the first example, Llama-2 attributes significant weights to Chunks 1 and 2, highlighting key grammatical mistakes.
+Conversely, it assigns a minor weight to Chunk 5 due to its imperfect modification.
+The second example shows Llama-2 attributing substantial weights to all chunks.
+Specifically, for Chunk 2, the hypothesis fails to remove the redundant “out," emphasizing the under-correction issue.
+Chunks 6 and 8 display excessive corrections, altering the sentence’s original intent and thus indicating considerable over-correction.
+Generally, Llama-2 tends to ascribe either very high or very low weights to modifications.
+We speculate it is due to the small scale of the LLM we adopt, impairing its ability to distinguish grammatical errors with varying levels.
+| Metric | EW | TS | ||
+| CLEME-dep-unw | 0.638 | 0.654 | 0.681 | 0.709 |
+| CLEME-ind-unw | 0.640 | 0.648 | 0.680 | 0.698 |
+| CLEME-dep-len | 0.700 | 0.665 | 0.691 | 0.742 |
+| CLEME-ind-len | 0.649 | 0.709 | 0.691 | 0.731 |
+| CLEME-dep-sim | 0.655 | 0.764 | 0.698 | 0.797 |
+| CLEME-ind-sim | 0.641 | 0.720 | 0.687 | 0.747 |
+| SentCLEME-dep-unw | 0.853 | 0.687 | 0.805 | 0.604 |
+| SentCLEME-ind-unw | 0.790 | 0.275 | 0.722 | 0.181 |
+| SentCLEME-dep-len | 0.876 | 0.824 | 0.844 | 0.808 |
+| SentCLEME-ind-len | 0.868 | 0.725 | 0.857 | 0.758 |
+| SentCLEME-dep-sim | 0.888 | 0.692 | 0.844 | 0.648 |
+| SentCLEME-ind-sim | 0.843 | 0.500 | 0.786 | 0.434 |
+C.3 Extra Results of CLEME with Similarity Weighting
+We additionally investigate the application of similarity-based weighting to CLEME Ye et al.
+(2023c) and present the results on CoNLL2014 in Table 10.
+We find that similarity-based weighting is superior to length-based weighting for corpus-level CLEME, while the trend is reversed for SentCLEME, and both are better than the unweighted setting.
+Moreover, it should be noted that no matter the weighting strategy employed, CLEME consistently underperforms compared to CLEME2.0.
+This is attributed to the fundamental disparities in design and scoring frameworks between the versions.
+CLEME2.0 was crafted to incorporate these sophisticated weighting techniques, allowing it to better distinguish between diverse error types and deliver a more thorough and refined performance assessment.
 | Metric | EW | TS | Avg. | ||
@@
 | SentCLEME2.0-sim-ind | 0.412 | 0.360 | 0.399 | 0.338 | 0.377 |
-| Dataset | Corpus-EW | Corpus-TS | Sentence-EW | Sentence-TS | ||||
-| CoNLL-2014 | 0.697 | 0.659 | 0.759 | 0.720 | 0.626 | 0.654 | 0.696 | 0.698 |
-| BN-10GEC | 0.732 | 0.764 | 0.796 | 0.813 | 0.638 | 0.637 | 0.708 | 0.698 |
-| E-Minimal | 0.709 | 0.786 | 0.779 | 0.819 | 0.642 | 0.692 | 0.715 | 0.747 |
-| E-Fluency | 0.760 | 0.786 | 0.831 | 0.841 | 0.642 | 0.665 | 0.720 | 0.714 |
-| NE-Minimal | 0.777 | 0.823 | 0.839 | 0.861 | 0.654 | 0.747 | 0.723 | 0.791 |
-| NE-Fluency | 0.823 | 0.692 | 0.849 | 0.709 | 0.664 | 0.791 | 0.742 | 0.830 |
-| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | Chunk 7 | Chunk 8 | Chunk 9 | |
-| Source | When we are | diagonosed out | with certain genetic | disease | , are we suppose to disclose | this result | to | our | relatives ? |
-| Ref. | When we are | diagnosed | with certain genetic | diseases | , are we suppose to disclose | this result | to | our | relatives ? |
-| Hyp. | When we are | diagnosed out (0.056) | with certain genetic | diseases (0.006) | , are we suppose to disclose | the results (0.019) | to | their (0.021) | relatives ? |
-| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | |
-| Source | Do one | who | suffered | from this disease keep it a secret | of infrom | their relatives ? |
-| Ref | Does one | who | suffers | from this disease keep it a secret | or inform | their relatives ? |
-| Hyp. | Do one (0.028) | who | suffer (0.011) | from this disease keep it a secret | to inform (0.094) | their relatives ? |
-5 Analysis
-5.1 Ablation Studies
-Isolated effect of each disentangled score.
-We perform ablation experiments on (Sent)CLEME2.0-dep to observe how each disentangled score performs.
-Since a system exhibiting lower error-correction, under-correction, and over-correction is more desirable, these scores are subtracted from 1.
-The results are presented in Table 3.
-Both hit-correction and under-correction scores display moderate correlations.
-Over-correction scores exhibit slight positive correlations at the corpus-level, but negligible negative correlations at the sentence level.
-Interestingly, error-correction scores manifest negative correlations.
-However, this does not imply that error-correction scores have no influence on the comprehensive score.
-In fact, the trade-off factor of error-correction scores is relatively significant.
-It is hypothesized that evaluations based solely on error-correction scores could unduly encourage systems that produce only highly confident edits, resulting in potential evaluation bias.
-Average correlations.
-To further compare different metrics from a global perspective, we report the average correlations obtained through the exhaustive enumeration of various parameter configurations.
-Specifically, all possible parameter combinations are attempted, with a step increment of 0.05.
-From Table 4, we observe that all the correlations are positive, regardless of the applied correction assumptions, evaluation levels, and weighting techniques.
-Comparing the unweighted and similarity-based weighted results, we conclude that similarity-based weighting significantly promotes human correlations, even on a global scale.
-Furthermore, corpus-level metrics tend to attain higher average results than sentence-level metrics; nonetheless, sentence-level metrics with optimal parameters surpass their corpus-level counterparts.
-This suggests that corpus-level metrics may exhibit enhanced robustness concerning parameter selection.
-5.2 Exploration of LLM
-The results of LLM-based edit weighting are shown in Table 5.
-The corpus-level results are quite satisfying and are comparable to those of most metrics such as PT-M2 and CLEME.
-However, the sentence-level outcomes are less satisfactory.
-This could be due to the fact that LLM assigns error-editing scores on a scale of 1 to 5, which is notably coarser when contrasted with the 0 to 1 continuous scoring scale.
-Sentence-level scores depend on the mean of the editing scores within a particular sentence.
-Consequently, even the slightest bias in the scores assigned by the LLM might lead to significant deviations in the sentence-level outcomes.
-Although the LLM has had some success, its performance still falls short when compared to the similarity-based weighting.
-This might be due to the scoring granularity of the 1 to 5 scale provided by the LLM not being sufficiently fine-tuned.
-In addition, the score heavily relies on the functionality of the LLM, which proves rather unstable.
-5.3 Case Study
-Table 6 presents examples of CLEME2.0.
-In the top group, chunk 2 achieves the highest score (0.056), highlighting its significant impact on the sentence.
-Although "diagnosed" was correctly amended, the omission of "out" persists, rendering the sentence still incorrect.
-Chunk 4 represents a hit-correction relating to the singular and plural forms in the source sentence and its low score indicates a minimal impact.
-Chunks 6 and 8 are types of over-correction.
-Chunk 6 does not change the original meaning, whereas chunk 8 introduces a more severe error due to an incorrect personal pronoun.
-In the second group, both chunks 3 and 5 exhibit error-corrections, with chunk 5 scoring higher than chunk 3.
-Chunk 3 involves issues of tense and singular-plural, while chunk 5 presents a more serious error that completely alters the meaning of the sentence.
-The weighting scores reflect the superiority of the method.
-For metrics that do not apply weightings, sorts of edits are uniformly assigned, which does not reflect the actual semantics.
-6 Conclusion
-This paper proposes CLEME2.0, an interpretable evaluation strategy, which are beneficial to reveal crucial characteristics of GEC systems.
-To address the limitations of traditional reference-based metrics in capturing semantic nuances, we enhance CLEME2.0 using two innovative edit weighting techniques: similarity-based and LLM-based weighting.
-Through comprehensive experiments and analyses, we validate the efficacy and reliability of our approach.
-We believe CLEME2.0 will provide a promising perspective on the task of grammatical error correction.
-Limitation
-Although CLEME2.0 can be extended to other languages, its effectiveness has not been tested in any language other than English.
-Furthermore, all reference sets utilized in our experiments are derived from the CoNLL-2014 shared task, which is a second language dataset.
-To demonstrate the robustness of our approaches, further experiments on evaluation datasets encompassing multiple languages and text domains are required.
-Finally, we strongly advocate for the construction of new GEC evaluation datasets to advance the development of NLP.
-Ethics Statement
-In this paper, we validate the effectiveness and robustness of our proposed approach using the CoNLL-2014, BN-10GEC, and SN-8GEC reference datasets.
-These datasets are sourced from publicly available resources on legitimate websites and do not contain any sensitive data.
-Additionally, all the baselines employed in our experiments are publicly accessible GEC metrics, and we have duly cited the respective authors.
-We confirm that all datasets and baselines utilized in our experiments are consistent with their intended purposes.
-Appendix A Details about GEC Meta-Evaluation
-A.1 Human Rankings
-GJG15 ranking.
-Grundkiewicz et al.
-(2015) propose the first large-scale human judgement dataset towards 12 participating systems of the CoNLL-2014 shared task.
-In this assessment, 8 native speaker are asked to rank the outputs of all the systems from best to worst.
-Two system ranking lists are generated using Expected Wins (EW) and TrueSkill (TS) respectively.
-Since all the involved systems are mostly classical systems such as statistical machine translation approaches Junczys-Dowmunt and Grundkiewicz (2014) and classifier-based approaches Rozovskaya et al.
-(2014), the dataset may not be a ideal test bed for meta-evaluation in the current time.
-SEEDA ranking.
-Kobayashi et al.
-(2024b) identify several limitations of GJG15 ranking dataset, and propose a new human ranking dataset called SEEDA.
-SEEDA consists of corrections with human ratings along two different granularities: edit-based and sentence-based, covering 12 state-of-the-art systems including large language models (LLMs), and two human corrections with different focuses.
-Three native English speakers participate in the annotation process.
-Similar to Grundkiewicz et al.
-(2015), the overall human rankings are derived from TrueSkill (TS) and Expected Wins (EW) based on pairwise judgments.
-| Item | CoNLL-2014 | BN-10GEC | E-Minimal | E-Fluency | NE-Minimal | NE-Fluency |
-| # Sents (Length) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) |
-| # Refs (Length) | 2,624 (22.8) | 13,120 (22.9) | 2,624 (23.2) | 2,624 (22.8) | 2,624 (23.0) | 2,624 (22.2) |
-| # Edits (Length) | 5,937 (1.0) | 36,677 (1.0) | 4,500 (1.0) | 8,373 (1.1) | 4,964 (0.9) | 11,033 (1.2) |
-| # Unchanged Chunks (Length) | 11,174 (4.8) | 93,496 (2.5) | 8,887 (6.3) | 12,823 (3.8) | 10,748 (5.1) | 14,086 (2.9) |
-| # Corrected/Dummy Chunks (Length) | 4,994 (1.3) | 26,948 (2.4) | 3,963 (1.2) | 6,305 (1.7) | 4,221 (1.2) | 6,892 (2.6) |
-Appendix B Statistics of Reference Datasets
-Table 7 presents the statistics of all the reference sets involved in our experiments.
-B.1 Baseline Metrics
-In our evaluation, we compare our method with the following reference-based baseline metrics, including corpus and sentence-level variants:
-M2 and SentM2 Dahlmeier and Ng (2012a) dynamically extract the hypothesis edits with the maximum overlap of gold annotations by utilizing the Levenshtein algorithm.
-GLEU and SentGLEU Napoles et al.
-(2015) are BLEU-like GEC metrics based on n-gram matching, rewarding hypothesis n-grams that align with the reference but not the source, while penalizing those aligning solely with the source.
-GLEU is the main metric in JFLEG, an English GEC dataset that highlights holistic fluency edits.
-ERRANT and SentERRANT Bryant et al.
-(2017) are the most mainstream GEC metrics, which are based on edit matching.
-They are able to extract edits more accurately, by utilizing the linguistically enhanced Damerau-Levenshtein algorithm.
-PT-M2 and SentPT-M2 Gong et al.
-(2022) leverage pre-trained language model (PLM) to evaluate GEC systems.
-The main idea is similar to M2 and ERRANT, but they can leverage the knowledge of pre-trained language models to score edits effectively.
-CLEME and SentCLEME Ye et al.
-(2023c) are proposed to provide unbiased scores for multi-reference evaluation.
-Besides, the authors introduce the correction independence assumption, so CLEME can perform based on either traditional correction dependence or correction independence assumptions.
-In addition, for the evaluation of SEEDA, we have additionally added the following evaluation methods in accordance with the evaluation methods reported in Kobayashi et al. (2024b):
-GoToScorer Gotou et al.
-(2020): takes into account the difficulty of error correction when calculating the evaluation score.
-The difficulty is calculated based on the number of systems that can correct errors.
-SOME Yoshimura et al. (2020b): optimizes human evaluation by fine-tuning BERT separately for criteria such as grammaticality, fluency, and meaning preservation.
-IMPARA Maeda et al. (2022): incorporates a quality assessment model fine-tuned using BERT parallel data and a similarity model that takes into account the effects of editing.
-Appendix C Detailed Results of Evaluation
-We list detailed evaluation results of CLEME2.0 on CoNLL-2014 in Table 8.
-Appendix D Experimental Details of LLM-based Edit Weighting
-Prompt:
-As a grammar correction evaluator, you are required to score the corrected editors for each grammatical error.
-We will give three lines, the first line is the original sentence given, the second line is the modification made to the editor, and the third line is the output form.
-The scoring range is 1-5. The larger the score, the more important the editor’s correction is. Correspondingly, the smaller the score, the less important the editor’s correction is.
-1 point indicates that this editor’s modification has almost no impact on the original sentence and is dispensable.
-2 points indicates that this editorial change has a slight impact.
-3 points indicates that this editor’s changes have a certain impact.
-4 points indicates that this editorial change is necessary.
-5 points indicates that this editing modification is very necessary and of high importance.
-The output format is a score of 1 to 5 points.
-Next, I will give you a sentence only with an edit. You need to rate each edit in sequence. The desired output is just a score, without any redundant explanation.
-Example Input:
-Sentence: Nowadays the technologies were improved a lot compared to the last century.
-Edit: were => have
-Output (1-5):
-Example Output:
-5
-Note that the output must be a number between 1 and 5. Here is the sample:
-Due to the strong semantic understanding capabilities of large language models (LLMs) Yang et al.
-(2020, 2022b, 2022a); Chai et al.
-(2024), recent work Sottana et al.
-(2023) has sparked interest in using LLMs for text evaluation, including the evaluation of grammatical error correction.
-Inspired by this, we utilize LLMs as weighted scorers to assess the importance of each edit.
-The template for the LLM is shown in Figure 3.
-For each edit, the constructed sentence contains only one grammatical error, while the other positions are correct.
-The second line shows the modification of that edit.
-The LLM is required to determine the necessity of the modified edit and output a score from 1 to 5.
-A higher score indicates a greater necessity for the edit modification.
-We do not inform the LLM of the specific types of edits; instead, we let the larger model evaluate the necessity of the modified edits.
-D.1 Hit-Correction Edits
-Scenario: The hypothesis and reference sentence are consistent.
-Focus: The significance of the transition from the source to the hypothesis sentence.
-Scoring: A higher score indicates that the edit from source to reference sentence carries substantial importance. Conversely, a lower score suggests that this transition is less crucial.
-D.2 Error-Correction Edits
-Scenario: The hypothesis and reference sentence are inconsistent.
-Focus: The significance of the transition from the hypothesis to the reference sentence.
-Scoring: A high score indicates a critical edit, suggesting significant inaccuracies in the hypothesis sentence.
-A low score implies that the modification is of minimal importance, indicating the hypothesis sentence is either correct or not substantially incorrect.
-D.3 Under-Correction Edits
-Scenario: The source and hypothesis sentence remain unchanged.
-Focus: The importance of modifications from the source to the reference sentence.
-Scoring: A high score implies a critical need for the edit, pointing to a severe under-correction.
-Conversely, a low score indicates that the edit is of lesser importance, suggesting a mild under-correction.
-D.4 Over-Correction Edits
-Scenario: The source is equivalent to the reference sentence, leading to two distinct situations:
-1.
-The reference is not an ideal sentence, and the hypothesis sentence is corrected but deemed overcorrected.
-2.
-The reference is optimal, necessitating no amendments, yet the hypothesis sentence introduces corrections.
-Evaluation:
-First Situation: Assess the importance () of the edit from the source to the hypothesis sentence.
-A higher W1 score indicates that the edit is crucial, suggesting imperfections in the reference sentence.
-Conversely, a lower score suggests that the edit is of minimal importance, rendering the hypothesis’s correction unnecessary.
-Second Situation: Examine the significance () of the edit from the hypothesis to the reference sentence.
-A higher score indicates that the edit is critical, denoting that the hypothesis’s correction was overly aggressive.
-A lower score implies the edit was unneeded, making the correction by the hypothesis irrelevant.
-Formula: The computation of over-correction score is defined as follow:
-This score can be either positive or negative. A higher over-correction score signals a less effective performance by the correction system.
-By systematically assessing the necessity and importance of different types of edits, we can better understand and improve the performance of grammatical error correction systems.
+C.4 Average Correlations.
+To analyze our method from a global viewpoint, we present the average correlations derived from the exhaustive enumeration of possible parameter configurations.
+We explore all potential parameter combinations with increments of 0.05.
+Table 11 shows that all correlations are positive, regardless of the correction assumptions, levels of evaluation, or weighting techniques used.
+By comparing results from unweighted and similarity-based weighted metrics, we determine that similarity-based weighting substantially enhances human correlation on a global level.
+Additionally, corpus-level metrics generally achieve higher average values compared to sentence-level metrics.
+However, sentence-level metrics with optimal parameters can outperform their corpus-level equivalents.
+This implies that corpus-level metrics might demonstrate greater robustness concerning parameter selection.
 | Metric | AMU | CAMB | CUUI | IITB | INPUT | IPN | NTHU | PKU | POST | RAC | SJTU | UFC | UMC | |
-| CLEME2.0-dependent | TP | 380 | 584 | 471 | 22 | 0 | 39 | 330 | 246 | 412 | 254 | 85 | 32 | 260 |
+| CLEME2.0-dep | TP | 380 | 584 | 471 | 22 | 0 | 39 | 330 | 246 | 412 | 254 | 85 | 32 | 260 |
 | sim | 9.20 | 12.66 | 7.58 | 0.39 | 0.00 | 0.77 | 5.79 | 6.69 | 8.80 | 6.68 | 1.50 | 0.42 | 5.29 | |
@@
 | sim | 34.25 | 28.45 | 36.21 | 78.39 | 48.24 | 83.10 | 46.53 | 36.15 | 34.48 | 38.00 | 56.28 | 51.27 | 39.60 | |
-| TN | 6298 | 6329 | 6224 | 6007 | 6308 | 5160 | 6274 | 6286 | 6428 | 6276 | 6313 | 5965 | 6355 | |
-| sim | 6237 | 6301 | 6190 | 5373 | 6226 | 5241 | 5973 | 6214 | 6382 | 6194 | 5902 | 6092 | 6273 | |
 | Hit | 0.188 | 0.271 | 0.220 | 0.010 | 0.00 | 0.013 | 0.163 | 0.126 | 0.198 | 0.127 | 0.046 | 0.015 | 0.132 | |
 | sim | 0.194 | 0.266 | 0.160 | 0.005 | 0.00 | 0.009 | 0.100 | 0.143 | 0.184 | 0.138 | 0.025 | 0.008 | 0.109 | |
-| Error | 0.137 | 0.194 | 0.145 | 0.016 | 0.00 | 0.048 | 0.150 | 0.130 | 0.152 | 0.130 | 0.042 | 0.006 | 0.125 | |
+| Wrong | 0.137 | 0.194 | 0.145 | 0.016 | 0.00 | 0.048 | 0.150 | 0.130 | 0.152 | 0.130 | 0.042 | 0.006 | 0.125 | |
 | sim | 0.086 | 0.138 | 0.078 | 0.009 | 0.00 | 0.052 | 0.101 | 0.0866 | 0.096 | 0.078 | 0.038 | 0.003 | 0.079 | |
@@
 | sim | 0.503 | 0.520 | 0.484 | 0.425 | 0.45 | 0.408 | 0.439 | 0.474 | 0.491 | 0.438 | 0.424 | 0.448 | 0.452 | |
-[…]
+| SentCLEME2.0-dep | TP | 376 | 580 | 467 | 22 | 0 | 39 | 327 | 244 | 409 | 251 | 84 | 32 | 259 |
 | sim | 9.14 | 12.63 | 7.52 | 0.39 | 0.00 | 0.76 | 5.72 | 6.65 | 8.75 | 6.59 | 1.48 | 0.42 | 5.23 | |
@@
 | sim | 43.65 | 35.92 | 45.22 | 57.46 | 58.31 | 54.69 | 46.92 | 46.02 | 43.09 | 46.05 | 55.32 | 58.35 | 48.02 | |
-| TN | 6058 | 6105 | 6004 | 6092 | 6108 | 6106 | 6002 | 6028 | 6195 | 6019 | 6081 | 6072 | 6102 | |
-| sim | 6052 | 6095 | 6009 | 6093 | 6106 | 6115 | 5995 | 6012 | 6203 | 6027 | 6079 | 6070 | 6115 | |
 | Hit | 0.136 | 0.210 | 0.163 | 0.008 | 0.00 | 0.013 | 0.119 | 0.088 | 0.142 | 0.089 | 0.032 | 0.012 | 0.091 | |
 | sim | 0.131 | 0.205 | 0.142 | 0.007 | 0.00 | 0.011 | 0.104 | 0.088 | 0.129 | 0.086 | 0.027 | 0.008 | 0.087 | |
-| Error | 0.080 | 0.129 | 0.090 | 0.005 | 0.00 | 0.038 | 0.095 | 0.076 | 0.088 | 0.071 | 0.023 | 0.002 | 0.070 | |
+| Wrong | 0.080 | 0.129 | 0.090 | 0.005 | 0.00 | 0.038 | 0.095 | 0.076 | 0.088 | 0.071 | 0.023 | 0.002 | 0.070 | |
 | sim | 0.063 | 0.102 | 0.066 | 0.004 | 0.00 | 0.033 | 0.079 | 0.059 | 0.070 | 0.051 | 0.020 | 0.001 | 0.059 | |
@@
 | sim | 0.502 | 0.520 | 0.504 | 0.467 | 0.466 | 0.449 | 0.479 | 0.481 | 0.494 | 0.484 | 0.467 | 0.469 | 0.479 | |
-| CLEME2.0-independent | TP | 388 | 596 | 487 | 22 | 0 | 39 | 338 | 248 | 420 | 255 | 85 | 32 | 262 |
+| CLEME2.0-ind | TP | 388 | 596 | 487 | 22 | 0 | 39 | 338 | 248 | 420 | 255 | 85 | 32 | 262 |
 | sim | 9.47 | 13.11 | 7.99 | 0.40 | 0.00 | 0.81 | 6.13 | 6.80 | 9.07 | 6.91 | 1.54 | 0.47 | 5.49 | |
@@
 | sim | 26.88 | 20.31 | 27.94 | 53.23 | 41.31 | 50.21 | 36.83 | 28.40 | 26.59 | 29.30 | 40.63 | 41.49 | 31.88 | |
-| TN | 6629 | 6701 | 6597 | 6567 | 6560 | 6664 | 6617 | 6611 | 6793 | 6628 | 6583 | 6546 | 6680 | |
 | Hit | 0.213 | 0.298 | 0.254 | 0.014 | 0.000 | 0.024 | 0.185 | 0.141 | 0.222 | 0.142 | 0.053 | 0.021 | 0.149 | |
 | sim | 0.222 | 0.298 | 0.193 | 0.007 | 0.000 | 0.015 | 0.117 | 0.160 | 0.212 | 0.165 | 0.035 | 0.011 | 0.126 | |
-| Error | 0.224 | 0.313 | 0.234 | 0.022 | 0.000 | 0.141 | 0.244 | 0.220 | 0.257 | 0.226 | 0.083 | 0.008 | 0.207 | |
+| Wrong | 0.224 | 0.313 | 0.234 | 0.022 | 0.000 | 0.141 | 0.244 | 0.220 | 0.257 | 0.226 | 0.083 | 0.008 | 0.207 | |
 | sim | 0.148 | 0.241 | 0.133 | 0.016 | 0.000 | 0.086 | 0.181 | 0.172 | 0.166 | 0.133 | 0.054 | 0.004 | 0.141 | |
@@
 | sim | 0.503 | 0.508 | 0.490 | 0.426 | 0.450 | 0.400 | 0.428 | 0.463 | 0.489 | 0.479 | 0.425 | 0.449 | 0.446 | |
-[…]
+| SentCLEME2.0-ind | TP-sim | 9.16 | 12.59 | 7.73 | 0.40 | 0.00 | 0.75 | 5.93 | 6.67 | 8.77 | 6.67 | 1.50 | 0.47 | 5.21 |
 | FP-sim | 15.83 | 29.93 | 15.62 | 1.76 | 0.00 | 12.58 | 24.30 | 14.17 | 18.94 | 12.00 | 6.84 | 0.27 | 17.76 | |
@@
 | sim | 0.154 | 0.240 | 0.174 | 0.009 | 0.000 | 0.014 | 0.125 | 0.100 | 0.155 | 0.103 | 0.033 | 0.012 | 0.102 | |
-| Error | 0.159 | 0.261 | 0.178 | 0.015 | 0.000 | 0.110 | 0.192 | 0.165 | 0.192 | 0.162 | 0.059 | 0.005 | 0.147 | |
+| Wrong | 0.159 | 0.261 | 0.178 | 0.015 | 0.000 | 0.110 | 0.192 | 0.165 | 0.192 | 0.162 | 0.059 | 0.005 | 0.147 | |
 | sim | 0.134 | 0.229 | 0.147 | 0.013 | 0.000 | 0.094 | 0.170 | 0.144 | 0.164 | 0.129 | 0.051 | 0.004 | 0.127 | |
@@
 | sim | 0.493 | 0.498 | 0.496 | 0.466 | 0.468 | 0.432 | 0.462 | 0.461 | 0.478 | 0.469 | 0.462 | 0.473 | 0.466 |
+C.5 Details Results on CoNLL-2014
+Table 12 presents a comprehensive evaluation of CLEME2.0 on CoNLL-2014 across all GEC systems.
+Our method offers a clear and quantitative examination of detailed features of GEC systems, which other automatic metrics cannot provide.
+For instance, the CAMB system attains the top hit-correction score of 0.271 for CLEME2.0-dep, which shows that about 27.1% of edits by the system are accurate.
+The wrong-correction score of 0.194 indicates that 19.4% of edits are correctly placed but incorrect, the under-correction score of 0.534 indicates that 53.4% of grammatical errors are overlooked by the system, and the over-correction score of 0.470 suggests that 47.0% of the edits are unnecessary.
+As a result, developers and researchers can pinpoint the aspects of their systems that require enhancement.
+Furthermore, users can select GEC systems that best meet their requirements.
+For instance, users might opt for a system with a minimal under-correction score in high-stakes situations, as they expect to detect every possible grammatical mistake even though the system might make some unnecessary edits.